### v1

# Chain-of-Influence: Tracing Interdependencies Across Time and Features in Clinical Predictive Modelings

Abstract
Modeling clinical time-series data is hampered by the challenge of capturing latent, time-varying dependencies among features. State-of-the-art approaches often rely on black-box mechanisms or simple aggregation, failing to explicitly model how the influence of one clinical variable propagates through others over time. We propose Chain-of-Influence (CoI), an interpretable deep learning framework that constructs an explicit, time-unfolded graph of feature interactions. CoI leverages a multi-level attention architecture: first, a temporal attention layer identifies critical time points in a patient’s record; second, a cross-feature attention layer models the directed influence from features at these time points to subsequent features. This design enables the tracing of influence pathways, providing a granular audit trail that shows how any feature at any time contributes to the final prediction, both directly and through its influence on other variables. We evaluate CoI on mortality and disease progression tasks using the MIMIC-IV dataset and a private chronic kidney disease cohort. Our framework significantly outperforms existing methods in predictive accuracy. More importantly, through case studies, we show that CoI can uncover clinically meaningful, patient-specific patterns of disease progression that are opaque to other models, offering unprecedented transparency into the temporal and cross-feature dependencies that inform clinical decision-making.
1 introduction
Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens National Kidney Foundation (2024); Murphy et al. (2019). The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging Lin et al. (2013); Li & Padman (2024).
Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities. However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care Caruana et al. (2015); Rudin (2019); Tonekaboni et al. (2019). This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME Ribeiro et al. (2016) and SHAP Lundberg & Lee (2017) toward more sophisticated attention-based approaches that provide intrinsic interpretability Choi et al. (2016; 2017); Bardhan (2024).
Despite these advances, current approaches to clinical prediction face fundamental limitations. Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships. Even sophisticated attention-based models like RETAIN Choi et al. (2016), while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
These chains of influence—where early clinical indicators propagate through interconnected pathways to affect later outcomes—represent a core challenge for prediction in both chronic and acute care settings. In the slow progression toward End-Stage Renal Disease (ESRD), a patient’s declining eGFR over several months may trigger medication adjustments that in turn affect subsequent laboratory values and hospitalization patterns. The timeline is compressed in an Intensive Care Unit (ICU), but the principle is the same: a sudden drop in blood pressure can necessitate a vasopressor, which then impacts heart rate and organ function within hours, directly influencing the immediate risk of mortality. Whether unfolding over years or hours, these complex temporal-feature interdependencies are what current models fail to capture explicitly.
Building on TFCAM, we introduce Chain-of-Influence (CoI), which adds hierarchical attention—temporal, feature-level, and cross-feature—and validates these choices with expanded technical details and broader experiments (Li et al., 2025b), enabling principled modeling and interpretation of temporal–feature interdependencies.
2 Related Work
Our work builds upon a rich history of attention-based models designed for clinical time-series data. We categorize prior work into three main areas: foundational attention architectures, models incorporating temporal and structural priors, and transformer-based approaches.
2.1 Foundational Attention Architectures
Attention mechanisms marked a turning point for interpretability in clinical deep learning. The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning Choi et al. (2016). By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions. Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies Ma et al. (2017). While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction. They do not explicitly capture the influence pathways—how a variable at one point in time may directly affect a different variable later on.
2.2 Time-Aware and Hierarchical Attention
A significant challenge in clinical data is its inherent irregularity and the varying importance of events over time. To address this, a second wave of research developed more sophisticated temporal attention models. For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time Zhang et al. (2019). Other models introduced domain-specific structural priors. StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time Gao et al. (2020). Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component Ma et al. (2020). These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
2.3 Self-Attention and Transformer-Based Approaches
The advent of the Transformer architecture, with its core self-attention mechanism, provided a powerful tool for capturing global, all-to-all interactions within a sequence Vaswani et al. (2017). This was quickly adapted to the clinical domain in models like BEHRT, which treated a patient’s record as a sequence of concepts and learned deep, contextualized representations Li et al. (2020). Subsequent work built hierarchical structures on this foundation. PAVE, for example, uses multi-layer self-attention to first identify significant clinical events and then aggregate them into higher-level, clinically meaningful event patterns Steinberg et al. (2021). More recently, XTSFormer introduced multi-scale temporal attention to distinguish between the influence of acute, short-term events and chronic, long-term patterns Zhang et al. (2024).
Self-attention excels at modeling a dense web of pairwise interactions. However, this strength can also be a weakness for interpretability in a clinical context. The resulting attention maps highlight a complex network of correlations but do not naturally expose the directed, sequential pathways of influence that characterize disease progression. They show which features co-occur or are correlated, but they do not provide a clear, traceable audit trail of how one event propagates its influence through other variables over time. Our proposed model, CoI, is designed specifically to bridge this gap, moving beyond identifying salient features to explicitly modeling and tracing these inter-feature, cross-time influence chains.
3 Methodology
In this section, we introduce the Chain-of-Influence (CoI) model, designed to capture and interpret both temporal dynamics and feature interactions in sequential clinical data. CoI integrates three layers of attention mechanisms—temporal attention, feature-level attention, and cross-feature attention—to provide both accurate predictions and rich interpretability.
3.1 Input Representation and Embedding
Given an input tensor $\mathbf{X}\in\mathbb{R}^{B\times T\times F}$ where $B$ is the batch size, $T$ is the number of time steps, and $F$ is the number of input features, the first stage is to project the input into a higher-dimensional embedding space. This is achieved via a learnable linear transformation:
$\mathbf{E}=\operatorname{Linear}(\mathbf{X}),\mathbf{E}\in\mathbb{R}^{B\times T\times D},$
where $D$ denotes the embedding dimension. To encode the temporal order, a fixed positional encoding is added to the embedded inputs, allowing the model to capture sequential information even before applying the attention mechanisms.
3.2 Temporal and Feature-Level Attention
Inspired by the two-level attention mechanism introduced in RETAIN Choi et al. (2016), we employ a dual-attention approach to separately extract temporal and feature-level importance. While RETAIN utilized GRU Cho et al. (2014) units, we implement two distinct bidirectional LSTM branches to better capture long-range dependencies and context from both past and future time steps. This bidirectional processing enables a more comprehensive understanding of temporal patterns, which is particularly crucial for detecting subtle interactions between features across different time points.
Temporal-Level Attention
A bidirectional LSTM is applied to the positionally encoded embeddings: $\mathbf{H}^{\text{temp }}=\operatorname{BiLSTM}(\mathbf{E})$ Subsequently, a linear layer projects the LSTM outputs to compute a scalar attention score $a_{t}$ for each time step: $\alpha_{t}=\operatorname{softmax}\left(\mathbf{w}^{\top}\mathbf{h}_{t}^{\text{temp }}\right)$ These temporal attention weights $\alpha_{t}$ reflect the contribution of each time step to the final prediction.
Feature-Level Attention
In parallel, another bidirectional LSTM processes the same embeddings to capture feature-specific dynamics: $\mathbf{H}^{\text{feat }}=\operatorname{BiLSTM}(\mathbf{E})$ A separate linear projection followed by a non-linear activation (e.g., tanh) is used to compute a vector of feature-level attention weights $\beta_{t}$ for each time step: $\beta_{t}=\tanh\left(\operatorname{Linear}\left(\mathbf{h}_{t}^{\text{feat }}\right)\right)$ Together, the temporal weights $\alpha_{t}$ and the feature-level weights $\beta_{t}$ yield a local contribution matrix $C$, where each element $C[t,i]$ is computed as:
$C[t,i]=\alpha_{t}\times\sum_{k}\left(\beta_{t}[k]\times W_{\mathrm{emb}}[k,i]\times x_{t}[i]\right),$
where $W_{\text{emb }}$ is the weight matrix of the initial embedding layer, and $x_{t}[i]$ is the $i$-th feature of the input at time $t$.
3.3 Cross-Feature Attention and Chained Influence
The cross-feature attention component leverages transformer-based layers to capture temporal dependencies and inter-feature interactions across time. In addition to learning complex patterns, these layers offer a pathway for interpretability by linking local contributions with cross-time interactions. Here, we detail how the cross-attention matrix $A$ is obtained, how the local contribution matrix $C$ is computed, and how they are combined into a chained influence measure.
Deriving the Cross-Attention Matrix AA
Within each transformer layer, multi-head self-attention is applied to the positionally encoded embeddings. For each sample, every layer outputs attention weights with dimensions $\text{num\_heads},T,T]$, where $T$ is the sequence length. These weights indicate the strength of interactions between time steps. To obtain a robust measure of these temporal relationships, we aggregate the attention weights by averaging over all heads and layers, resulting in the cross-attention matrix:
$A\in\mathbb{R}^{T\times T}.$
In our notation, each element $A[t,t^{\prime}]$ represents the attention weight from time step $t^{\prime}$ to time step $t$ (with $t<t^{\prime}$), quantifying how much the information at time step $t$ contributes to the representation at time step $t^{\prime}$.
The matrix $A$ is computed by aggregating the attention weights from all transformer layers and heads. In particular, if $A^{(l,h)}$ denotes the attention weight matrix from the $h$-th head in the $l$-th layer, then we define:
$A[t,t^{\prime}]=\frac{1}{L\cdot H}\sum_{l=1}^{L}\sum_{h=1}^{H}A^{(l,h)}[t,t^{\prime}],$
where $L$ is the number of transformer layers and $H$ is the number of attention heads per layer. This aggregated matrix $A$ captures how information is transmitted from earlier to later time steps. Specifically, a higher value of $A[t,t^{\prime}]$ indicates that the representation at time $t^{\prime}$ heavily incorporates information from time $t$, thereby revealing long-range dependencies and offering insights into the model’s internal temporal dynamics.
Computing the Local Contribution Matrix $C$
As detailed in Section 2.1.2, the local contribution matrix $C$ encapsulates the importance of each feature at every time step by integrating the temporal attention weights ($\alpha_{t}$) and the feature-level attention vectors ($\mathbf{b}_{t}$). Combined with the embedding layer’s weight matrix $W_{\mathrm{emb}}$ and the raw input $x_{t}$, the contribution of feature $i$ at time $t$ is computed as:
$C[t,i]=\alpha_{t}\times\sum_{k}\left(b_{t}[k]\times W_{\mathrm{emb}}[k,i]\times x_{t}[i]\right)$
In this expression, the scalar $\alpha_{t}$ captures the overall relevance of time step $t$, while the vector $\mathbf{b}_{t}$ distributes attention across the features at that time. The summation over $k$ aggregates the contribution across the embedding dimensions, effectively translating the feature-level weights into the original input space. Thus, $C[t,i]$ serves as a composite measure that reflects not only the direct impact of feature $i$ at time $t$ but also the modulation of this impact by the importance of the time step and the relative emphasis on the specific feature. It provides a fine-grained interpretation of the direct and modulated impact of feature $i$ at time $t$.
Quantifying Temporal-Feature Cross Influence
Having established the local contribution matrix $C$ and the temporal connectivity matrix $A$, we then seek to quantify explicitly how the importance of a specific feature at an earlier time step affects another feature at a subsequent time step. Specifically, our goal is to detect and measure how feature $i$ at time $t$ influences feature $j$ at time $t^{\prime}$(with $t<t^{\prime}$).
To integrate the local feature importance with temporal relationships, we define a chained influence measure that quantifies the effect of an earlier feature on a later one:
$\text{I}(t,i;t^{\prime},j)=C[t,i]\times A[t,t^{\prime}]\times C[t^{\prime},j],$
where $C[t,i]$ represents the local contribution of feature $i$ at time $t$, $C[t^{\prime},j]$ represents the local contribution of feature $j$ at time $t^{\prime}$, and $A[t,t^{\prime}]$ denotes the attention weight that quantifies the extent to which information from time $t$ is incorporated into the representation at time $t^{\prime}$. This formulation effectively captures the dynamic interplay between individual feature contributions and their propagation through time, offering a detailed and interpretable account of how early features influence later ones within the CoI framework.
3.4 Implementation Optimizations
To achieve high computational efficiency and robust performance, we introduce two key implementation optimizations:
Vectorized Computation of Temporal-Feature Cross Influence: To compute the chained influence tensor efficiently, we leverage array stacking and the Einstein summation convention via np.einsum. This vectorized approach replaces nested loops with concise batch-level operations, dramatically improving computational efficiency and scalability. Specifically, the influence tensor is computed as:
$\mathrm{I}[n,t,u,i,j]=A[n,t^{\prime},t]\times C[n,t,i]\times C[n,t^{\prime},j],$
where $n$ indexes samples, $t$ and $t^{\prime}$ index time steps (with $t<t^{\prime}$), and $i$ and $j$ index features.
Dynamic Tanh (DyT) Integration: We replace traditional normalization layers with Dynamic Tanh (DyT) activation function Zhu et al. (2025), which dynamically adapts the non-linearity based on the input distribution:
$\text{DyT}(x)=\alpha\cdot\tanh(\beta\cdot x+\gamma),$
where $\alpha$, $\beta$, and $\gamma$ are learnable parameters.
4 Experimental Setup
4.1 Datasets
We evaluate our Chain-of-Influence framework on two complementary clinical datasets representing distinct prediction scenarios: chronic disease progression and acute care outcomes.
4.1.1 Chronic Kidney Disease (CKD) Dataset
Our primary evaluation utilizes a proprietary longitudinal CKD dataset capturing gradual kidney function deterioration over extended periods, making it ideal for evaluating temporal-feature interdependencies in chronic disease progression.
The cohort comprises 1,422 CKD patients monitored over 24 months at 3-month intervals (8 time points), with 38 clinical features spanning demographics, comorbidities, laboratory biomarkers, and healthcare utilization. The prediction target is progression to End-Stage Renal Disease (ESRD), occurring in 6.0% of patients—a critical endpoint requiring dialysis or transplantation.
4.1.2 MIMIC-IV Dataset
To demonstrate generalizability, we evaluate on MIMIC-IV v3.1 Johnson et al. (2023), a public critical care database. From 94,458 ICU stays, we constructed a cohort of 65,366 patients using their first ICU stay, extracting hourly measurements over 48 hours across 15 clinical features (vital signs, laboratory tests, demographics).
Unlike the CKD dataset’s monthly resolution, MIMIC-IV provides hourly temporal data capturing rapid physiological changes characteristic of critical illness. The prediction target is in-hospital mortality (10.8% prevalence).
These complementary datasets enable evaluation across different temporal scales (hours vs. months) and clinical contexts (acute vs. chronic care). Comprehensive dataset characteristics, feature specifications, and preprocessing details are provided in Appendix A.
4.2 Baselines
We compare Chain-of-Influence (CoI) against three representative baselines to benchmark its performance. We include a standard BiLSTM as a strong but non-interpretable recurrent baseline that processes sequences bidirectionally. To compare against models with explicit temporal attention, we use RETAIN, the canonical interpretable baseline with a dual-level attention mechanism. Finally, to assess CoI against a different attention paradigm with strong clinical priors, we use StageNet, which leverages channel-wise attention to re-weight features based on disease stage. This selection provides a comprehensive benchmark against a standard recurrent model (BiLSTM), a model with explicit temporal attention (RETAIN), and one focused on channel-wise attention (StageNet).
4.3 Evaluation Metrics
We evaluate model performance using standard classification metrics. The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) Hanley & McNeil (1982); Fawcett (2006), which measures the model’s discriminative ability across all classification thresholds. We also report the F1-Score Van Rijsbergen (1979), the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets. Additionally, we assess overall prediction accuracy Sokolova & Lapalme (2009), which captures the proportion of correctly classified instances across both classes. Finally, we examine precision Van Rijsbergen (1979), which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions Powers (2011).
5 Results & Analysis
5.1 Predictive Performance
Model CKD Dataset MIMIC-IV Dataset
AUROC F1 Acc Prec Recall AUROC F1 Acc Prec Recall
Bi-LSTM 0.910 0.616 0.900 0.700 0.550 0.700 0.267 0.890 0.400 0.200
RETAIN 0.930 0.671 0.920 0.760 0.600 0.944 0.667 0.900 0.500 1.000
StageNet 0.940 0.685 0.930 0.780 0.610 0.920 0.768 0.930 0.700 0.850
CoI 0.960 0.721 0.950 0.820 0.660 0.950 0.865 0.950 0.850 0.880
Table 1: Performance comparison across datasets.
The comparative evaluation demonstrates the robust performance of Col across two distinct clinical prediction tasks. On the chronic CKD dataset, Col establishes clear superiority, outperforming the next-best model, StageNet, across all metrics. It achieves a notable 5.3% improvement in F1-score ( 0.721 vs. 0.685 ) and a $5.1\%$ increase in precision ( 0.820 vs. 0.780 ). This highlights Col’s enhanced ability to model the slow, interconnected progression of chronic disease over long timeframes.
On the acute care MIMIC-IV dataset, Col again delivers state-of-the-art performance, leading in AUROC, F1-score, Accuracy, and Precision. The results present a clinically significant nuance when comparing Col to RETAIN. While RETAIN achieves perfect recall (1.000), it does so with very low precision (0.500), indicating a high rate of false positives. In a critical care setting, such a high false alarm rate can lead to unnecessary interventions and alarm fatigue. In contrast, Col achieves a much stronger and more balanced profile, with high precision ( 0.850 ) and recall (0.880), culminating in a far superior F1-score (0.865 vs. RETAIN’s 0.667).
The performance of the baselines reveals key architectural insights. The standard Bi-LSTM performs adequately on the CKD data but struggles significantly with the complex, high-frequency interactions in the MIMIC-IV dataset ( $\mathrm{F}1=0.267$ ), showcasing the limitations of standard recurrent models in acute care. RETAIN’s behavior suggests its dual-attention mechanism is highly sensitive in the acute setting, adept at catching all true cases but at the cost of clinical utility due to low precision. StageNet proves to be a strong and balanced baseline on both datasets, making Col’s performance gains over it particularly meaningful.
Ultimately, Col’s strength lies in its architectural robustness. Its ability to achieve top-tier performance in both a sparse, long-term chronic setting and a dense, short-term acute care setting demonstrates that explicitly modeling feature influence chains captures generalizable patterns of deterioration. This suggests our framework has broad applicability across diverse clinical domains.
5.2 Interpretability Analysis
Due to space constraints, we present Chain-of-Influence’s interpretability capabilities using the CKD dataset, whose complex temporal dynamics and extended observation periods provide an ideal framework for demonstrating influence chain analysis.
5.2.1 Temporal Attention Patterns
Figure 1: Temporal attention comparison between RETAIN and CoI models. CoI (red) shows pronounced attention focusing on final time steps ($t_{6}$-$t_{7}$) with peak weights reaching 0.29, while RETAIN (blue) maintains more uniform distribution with maximum attention of 0.22 at $t_{7}$. The U-shaped pattern of CoI reflects clinical understanding that both baseline measurements and pre-progression periods are most informative for ESRD prediction.
Figure 1 presents a striking comparison of temporal attention patterns between CoI and RETAIN across the 8-time-point CKD progression timeline. The visualization reveals fundamentally different attention strategies: while RETAIN exhibits relatively uniform attention distribution with gradual increases toward recent time points, CoI demonstrates dramatic temporal focusing with a distinct U-shaped pattern that culminates in peak attention at the final time steps.
The temporal attention analysis reveals several clinically significant patterns. Most notably, CoI exhibits minimal attention weights (approximately 0.08) during the middle monitoring periods ($t_{1}$-$t_{3}$), followed by a dramatic escalation beginning at $t_{4}$ and reaching maximum attention at $t_{6}$-$t_{7}$ (0.29). This attention pattern aligns remarkably well with clinical understanding of CKD progression, where the final monitoring periods before ESRD onset contain the most prognostically valuable information. In contrast, RETAIN’s attention increases more gradually from 0.09 at $t_{1}$ to 0.22 at $t_{7}$, failing to distinguish between routine monitoring periods and critical pre-progression phases.
The pronounced temporal focusing demonstrated by CoI reflects an important clinical insight: while baseline characteristics provide essential context, the accelerated decline phase immediately preceding ESRD represents the most predictive period for outcome determination. This pattern suggests that CoI successfully learns to identify the temporal windows when cascading physiological changes become irreversible, enabling more precise risk stratification than models with uniform temporal weighting.
5.2.2 Feature Importance Analysis
(a) Feature importance scores for RETAIN model on CKD dataset.
(b) Feature importance scores for CoI model on CKD dataset.
Figure 2: Comparison of feature attributions on CKD. Normalized feature-importance profiles for (a) RETAIN and (b) CoI on the CKD dataset, shown side-by-side to compare salient predictors and attribution sparsity. Values are on the same scale to enable direct visual comparison; higher bars indicate greater contribution to the model’s decision.
While traditional feature importance rankings provide valuable insights, our Chain-of-Influence approach reveals the limitations of static importance scores in capturing temporal dependencies. Figure 2 presents detailed feature importance comparisons between RETAIN and CoI, showing both models consistently identify diabetes as the primary risk factor, corroborating established clinical evidence Burckhardt et al. (2017); Li et al. (2025a). Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels consistently rank among top features across both models.
However, static importance rankings obscure the temporal dynamics that CoI’s influence chains reveal. While CoI appropriately ranks eGFR higher ($6^{th}$) than RETAIN ($14^{th}$), neither model captures how eGFR’s influence propagates through temporal cascades to drive healthcare utilization patterns. Similarly, traditional rankings show diabetes as most important but fail to reveal how its influence chains through cardiovascular complications to accelerate kidney function decline over specific time horizons.
The Chain-of-Influence paradigm transcends these limitations by mapping how feature importance evolves temporally. Rather than asking which features are most important, CoI reveals when features become influential and how their effects cascade through clinical pathways. This temporal perspective explains why features like age demonstrate broad systemic influence across multiple chains while race shows targeted effects within specific progression pathways, providing clinicians with actionable insights beyond static importance scores.
5.2.3 Chain-of-Influence Analysis
Figure 3: Chain-of-Influence network visualization for CKD progression prediction. Nodes represent clinical features at specific time points (e.g., eGFR_t-5 indicates eGFR at 5 time steps before prediction). Directed edges capture learned temporal dependencies between features, with red edges indicating risk-enhancing influences and blue edges representing protective relationships. The network reveals complex cascading pathways from early cardiovascular events through kidney function decline to healthcare utilization escalation, demonstrating CoI’s ability to capture clinically meaningful temporal-feature interactions that drive disease progression.
Figure 3 reveals CoI’s breakthrough capability to map chains of influence across the 24-month CKD progression timeline. The network exposes how clinical deterioration propagates through interconnected pathways: cardiovascular complications (CHF_t-6, CD_t-5) trigger cascading effects through kidney function decline (eGFR_t-5, eGFR_t-6) to advanced disease stages (S4_t-5, S5_t-6), ultimately culminating in healthcare system escalation (n_claims_I_t-4, net_exp_DR_t-6). This temporal chaining quantifies the established cardio-renal syndrome with unprecedented precision, transforming clinical intuition into measurable influence pathways.
Critically, CoI identifies persistent influence chains that capture the irreversible nature of advanced clinical conditions. The S5_t-5 → S5_t-6 → S5_t-7 progression exemplifies how CKD Stage 5 creates self-reinforcing temporal chains, where each occurrence dramatically amplifies subsequent disease burden and progression risk. Similarly, CHF demonstrates persistent chaining patterns (CHF_t-6 → CHF_t-4 → CHF_t-1), reflecting the chronic, progressive nature of heart failure that compounds kidney disease severity. These persistent chains transcend simple temporal adjacency—they represent traced pathways of irreversible pathophysiological damage that fundamentally alter patient trajectories, enabling CoI to distinguish between transient clinical events and permanent disease state transitions.
The influence chains demonstrate remarkable bidirectional complexity, particularly evident in healthcare utilization patterns. CoI discovers that increased outpatient visits (n_claims_O_t-7) and professional consultations (n_claims_P_t-6) form feedback loops—simultaneously reflecting disease progression while driving improved outcomes through enhanced monitoring. This bidirectional influence challenges traditional unidirectional prediction models and reveals how clinical interventions reshape disease trajectories through temporal chains of care optimization.
CoI’s selective demographic integration further validates the chain-of-influence paradigm. Age propagates broad systemic effects across multiple clinical pathways, while gender and race demonstrate targeted, condition-specific influence patterns within discrete chains. The red-blue edge visualization immediately distinguishes risk-enhancing from protective influence chains, providing clinicians with the first quantitative map of CKD’s temporal complexity. This represents a paradigm shift from static risk scoring to dynamic influence tracking, enabling precise intervention timing based on evolving chain dynamics.
6 Limitations
We acknowledge key limitations: the learned influence chains represent powerful statistical associations, not formal causal pathways, and our aggregation of attention weights may obscure finer-grained component roles. Furthermore, the framework’s $O\left(T^{2}\cdot F^{2}\right)$ computational complexity presents a scalability challenge for datasets with extremely long time series or numerous features. Future work could address these by integrating causal discovery methods and exploring sparse approximation techniques.
7 Conclusion
In summary, the Chain-of-Influence framework advances the state-of-the-art in clinical predictive modeling by bridging the gap between performance and interpretability. Our approach explicitly models and visualizes the temporal and cross-feature dependencies that underlie complex clinical outcomes, achieving robust performance gains across heterogeneous datasets while providing clinicians with actionable, transparent explanations. CoI’s dual attention and influence tracing mechanism transforms black-box deep learning predictions into interpretable, dynamic influence maps that mirror clinical reasoning and inform optimal intervention strategies.
Looking forward, the flexibility of the Chain-of-Influence paradigm suggests broad applicability across diverse temporal and multi-modal clinical scenarios. Future work will explore extending the framework to handle richer data modalities, finer-grained temporal granularity, and integration with real-time clinical decision systems. By providing a transparent and generalizable approach to understanding the interplay of clinical variables over time, CoI represents a critical step toward safer, more accountable, and more effective deployment of AI in healthcare.
Appendix A Dataset Characteristics
A.1 Chronic Kidney Disease Dataset Statistics
Cohort Overview and Comparative Analysis
The Chronic Kidney Disease (CKD) dataset represents a comprehensive longitudinal study of 1,422 patients with progressive kidney disease, followed over 24 months. Table 2 presents a comparative analysis between patients who progressed to ESRD versus those who did not, providing critical insights into risk factors and disease progression patterns.
Table 2: CKD Dataset: Comprehensive Patient Characteristics by ESRD Progression Status
Characteristic ESRD Progression No ESRD Progression P-value
(n=86, 6.0%) (n=1,336, 94.0%)
Demographics
Age (years) $69.13\pm 12.37$ $72.04\pm 11.25$ <0.001<0.001
Female $40$ (46.5%) $721$ (54.0%) $0.215$
Race <0.001<0.001
   White $70$ (81.4%) $1,242$ (93.0%)
   African American $12$ (14.0%) $60$ (4.5%)
   Other 44 (4.6%) $34$ (2.5%)
BMI (kg/m²) $28.40\pm 5.32$ $26.40\pm 6.20$ <0.001<0.001
Comorbidities
Diabetes $63$ (73.3%) $788$ (59.0%) $0.009$
Hypertension $85$ (99.0%) $1,323$ (99.0%) $0.863$
Cardiovascular Disease $10$ (17.2%) $177$ (18.2%) $0.990$
Anemia $55$ (64.0%) $828$ (62.0%) $0.714$
Metabolic Acidosis $22$ (25.6%) $240$ (18.0%) $0.077$
Proteinuria $11$ (12.8%) $227$ (17.0%) $0.312$
Secondary Hyperparathyroidism $28$ (32.6%) $240$ (18.0%) <0.001<0.001
Phosphatemia 44 (4.7%) $40$ (3.0%) $0.390$
Atherosclerosis 66 (9.4%) $149$ (14.6%) $0.320$
Heart Failure 66 (7.0%) $120$ (9.0%) $0.526$
Stroke 11 (1.2%) $40$ (3.0%) $0.506$
Conduction & Dysrhythmias 44 (4.7%) $214$ (16.0%) $0.005$
Myocardial Infarction $31$ (51.7%) $316$ (32.4%) $0.003$
Fluid & Electrolyte Disorders 99 (17.6%) $122$ (13.3%) $0.503$
Metabolic Disorders 55 (9.4%) $60$ (6.5%) $0.581$
Nutritional Disorders 66 (10.2%) $106$ (11.6%) $0.900$
CKD Stage 4 $47$ (54.7%) $298$ (22.3%) <0.001<0.001
CKD Stage 5 $42$ (48.8%) $277$ (20.7%) <0.001<0.001
Feature Categories and Descriptions
The 38 clinical features in the CKD dataset are systematically organized across four major categories, each capturing distinct aspects of patient health and disease progression.
Table 3: CKD Dataset: Complete Feature Specifications
Category Feature Description
Demographics Age Patient age at baseline (years)
Gender Binary indicator (Male=0, Female=1)
Race Categorical encoding (White, Black, Hispanic, Other)
BMI Body Mass Index (kg/m²)
Comorbidities Diabetes Type 1 or Type 2 diabetes mellitus diagnosis
Htn Hypertension (systolic $\geq$140 or diastolic $\geq$90 mmHg)
Cvd Cardiovascular disease (coronary, cerebrovascular, peripheral)
Anemia Hemoglobin <<13 g/dL (men) or <<12 g/dL (women)
MA Mineral and bone disorders
Prot Proteinuria (>>300 mg/day or >>300 mg/g creatinine)
SH Secondary hyperparathyroidism
Phos Phosphorus disorders (hyperphosphatemia)
Athsc Atherosclerosis
CHF Congestive heart failure
Stroke Cerebrovascular accident history
CD Coronary disease
MI Myocardial infarction history
FE Iron deficiency
MD Metabolic disorders
ND Nutritional deficiency
S4 CKD Stage 4 (eGFR 15-29 mL/min/1.73m²)
S5 CKD Stage 5 (eGFR <<15 mL/min/1.73m²)
Lab Biomarkers Serum_Calcium Serum calcium levels (mg/dL)
eGFR Estimated glomerular filtration rate (mL/min/1.73m²)
Phosphorus Serum phosphorus levels (mg/dL)
Intact_PTH Intact parathyroid hormone (pg/mL)
Hemoglobin Blood hemoglobin concentration (g/dL)
UACR Urine albumin-to-creatinine ratio (mg/g)
Bicarbonate Serum bicarbonate levels (mEq/L)
Healthcare Utilization n_claims_DR Number of durable medical equipment/drug claims
n_claims_I Number of inpatient claims
n_claims_O Number of outpatient claims
n_claims_P Number of physician service claims
net_exp_DR Net expenditure for DME/drugs (\$)
net_exp_I Net expenditure for inpatient services (\$)
net_exp_O Net expenditure for outpatient services (\$)
net_exp_P Net expenditure for physician services (\$)
Table 4: CKD Dataset: Laboratory Values and Healthcare Utilization by ESRD Progression Status
Feature ESRD Progression No ESRD Progression P-value
(n=86) (n=1,336)
Laboratory Biomarkers
eGFR (mL/min/1.73m²) $17.21\pm 5.46$ $22.78\pm 5.66$ <0.001<0.001
Hemoglobin (g/dL) $12.15\pm 2.19$ $14.25\pm 1.80$ <0.001<0.001
Bicarbonate (mEq/L) $22.90\pm 6.36$ $25.30\pm 4.22$ $0.001$
Serum Calcium (mg/dL) $9.39\pm 3.62$ $10.21\pm 2.86$ $0.042$
Phosphorus (mg/dL) $3.61\pm 0.87$ $3.52\pm 0.72$ $0.350$
Intact PTH (pg/mL) $78.66\pm 40.23$ $62.72\pm 37.32$ $0.001$
Healthcare Utilization Claims
Pharmacy Claims (count) $120\pm 94$ $109\pm 86$ $0.293$
Inpatient Claims (count) $3.85\pm 3.41$ $3.74\pm 3.62$ $0.773$
Outpatient Claims (count) $7.78\pm 24.75 22$$2.07\pm 19.13$ $0.039$
Professional Claims (count) $105.37\pm 77.56$ $87.43\pm 68.02$ $0.039$
Healthcare Expenditures (\$)
Pharmacy Expenditure $12,053\pm 11,596$ $10,440\pm 20,662$ $0.242$
Inpatient Expenditure $33,909\pm 53,540$ $29,440\pm 32,541$ $0.446$
Outpatient Expenditure $9,354\pm 17,522$ $8,554\pm 17,492$ $0.682$
Professional Expenditure $5,512\pm 18,657 11$$1,640\pm 12,748$ $0.061$
Statistical Analysis and Clinical Significance
Cohort Characteristics: The complete cohort comprises 1,422 CKD patients with 24-month follow-up, yielding 8 temporal observations per patient (3-month intervals). ESRD progression occurred in 86 patients (6.0%), representing a clinically realistic progression rate for this advanced CKD population.
Significant Risk Factors (p << 0.05): Statistical analysis revealed several key differentiators between ESRD progressors and non-progressors:
Demographic Factors:
  • Younger age (69.1 vs 72.0 years) - counterintuitive finding suggesting more aggressive disease in younger patients
  • Higher BMI (28.4 vs 26.4 kg/m²) - indicating metabolic burden
  • Higher proportion of African Americans (14.0% vs 4.5%) - known ethnic disparity in CKD progression
Comorbidity Profile:
  • Higher diabetes prevalence (73.3% vs 59.0%) - primary driver of CKD progression
  • Increased secondary hyperparathyroidism (32.6% vs 18.0%) - marker of mineral bone disorder
  • More myocardial infarction (51.7% vs 32.4%) - cardiovascular complications
  • Paradoxically lower conduction disorders (4.7% vs 16.0%) - may reflect survival bias
  • Higher CKD stage 4 (54.7% vs 22.3%) and stage 5 (48.8% vs 20.7%) prevalence
Laboratory Markers:
  • Significantly lower eGFR (17.2 vs 22.8 mL/min/1.73m²) - expected primary marker
  • Lower hemoglobin (12.2 vs 14.3 g/dL) - indicating CKD-related anemia
  • Lower bicarbonate (22.9 vs 25.3 mEq/L) - metabolic acidosis marker
  • Higher intact PTH (78.7 vs 62.7 pg/mL) - secondary hyperparathyroidism
  • Lower serum calcium (9.4 vs 10.2 mg/dL) - mineral bone disorder
Healthcare Utilization:
  • Higher outpatient claims (27.8 vs 22.1) - increased monitoring needs
  • Higher professional claims (105.4 vs 87.4) - specialist consultations
A.2 MIMIC-IV Dataset Characteristics
Cohort Construction and Selection Criteria
The MIMIC-IV cohort was constructed through a systematic pipeline designed to ensure data quality and clinical relevance for mortality prediction tasks.
Initial Database Statistics:
  • Total ICU stays in MIMIC-IV: 94,458
  • Unique patients: 65,366
  • Multiple ICU stays per patient: 29,092 additional stays
Cohort Construction: Our preprocessing pipeline selects the first ICU stay per patient to ensure independence of observations, resulting in exactly 65,366 patients. This approach eliminates potential confounding from repeated admissions while maintaining a substantial sample size for robust model training and evaluation.
Final Cohort: 65,366 patients meeting all inclusion criteria.
Feature Statistics
Table LABEL:tab:mimic_detailed_stats presents detailed statistics for all MIMIC-IV features, including percentile distributions to capture the full range of physiological values encountered in critical care.
Table 5: MIMIC-IV Dataset: Comprehensive Feature Statistics (n=65,366 patients)
Feature Mean ± SD Median [IQR] Missing (%)
Demographics
Age (years) $4.5\pm 17.1 66$.0[54.0,78.0]66.0[54.0,78.0] $0.0$
Gender (Female %) $43.8\%$ - $0.0$
Vital Signs
Heart Rate (bpm) $84.1\pm 18.9^{*}$ $82.0[71.0,95.0]$ $27.5$
Systolic BP (mmHg) $19.9\pm 25.8^{*} 11$7.0[104.0,133.0]117.0[104.0,133.0] $51.6$
Diastolic BP (mmHg) $67.4\pm 15.8^{*}$ $64.0[55.0,75.0]$ $51.6$
Mean BP (mmHg) $87.6\pm 16.9^{*}$ $78.0[69.0,89.0]$ $51.6$
Respiratory Rate (rpm) $22.7\pm 6.8^{*}$ $19.0[15.0,22.0]$ $28.4$
Temperature (°C) $37.0\pm 0.8$ $36.8[36.6,37.2]$ $81.3$
SpO2 (%) $96.8\pm 3.2^{*}$ $97.0[95.0,99.0]$ $28.9$
Laboratory Tests
Creatinine (mg/dL) $1.4\pm 1.5$ $1.0[0.7,1.5]$ $93.1$
Glucose (mg/dL) $141.1\pm 69.7$ $125.0[103.0,157.0]$ $93.5$
Sodium (mEq/L) $138.2\pm 5.6$ $138.0[135.0,141.0]$ $92.9$
Potassium (mEq/L) $4.2\pm 0.7$ $4.1[3.8,4.5]$ $92.8$
Hematocrit (%) $31.1\pm 6.0$ $30.5[26.7,35.1]$ $92.3$
WBC (K/µL) $12.3\pm 10.7$ $10.7[7.8,14.7]$ $93.3$
Note: indicates features with outliers present in raw data that require preprocessing. Standard deviations shown are after outlier detection but before winsorization.
Outlier Detection and Treatment: Physiologically implausible values are identified using clinical knowledge-based thresholds and statistical methods:
  • Heart Rate: <<20 or >>220 bpm
  • Blood Pressure: Systolic <<50 or >>250 mmHg, Diastolic <<20 or >>150 mmHg
  • Temperature: <<32°C or >>42°C
  • SpO2: <<70% or >>100%
  • Laboratory values: Beyond 99.5th percentile of physiologically reasonable ranges
Outliers are winsorized to the 1st and 99th percentiles to preserve distributional properties while reducing the impact of measurement errors.
Normalization: All continuous features are standardized using z-score normalization: $z=\frac{x-\mu}{\sigma}$, where $\mu$ and $\sigma$ are computed on the training set only to prevent data leakage.
Cohort Characteristics and Outcomes
Table 6: MIMIC-IV Cohort: Clinical Characteristics and Outcomes
Characteristic Value Percentage
Total Patients 65,366 -
Age Distribution
   Mean Age $64.5\pm 17.1$ years -
   Age Range 18-103 years -
Gender
   Male 36,720 56.2%
   Female 28,646 43.8%
Temporal Structure
   Observation Window 48 hours Hourly resolution
   Total Observations 3,137,568 (65,366 × 48)
   Features per Patient 15 7 vitals + 6 labs + 2 demographics
Clinical Outcomes
Hospital Mortality 7,086 10.8%
Survivors 58,280 89.2%
Data Quality Characteristics
Vital Signs Missing Data 27.5-81.3% Variable by feature type
Laboratory Tests Missing Data 92.3-93.5% Expected for ICU setting
Demographics Missing Data 0.0% Complete for all patients
Appendix B Data Processing Implementation Details
B.1 MIMIC-IV Preprocessing Pipeline
The preprocessing pipeline for MIMIC-IV data involves several computational steps designed to handle the scale and complexity of the raw database:
Computational Requirements:
  • Memory usage: ~32GB RAM for full dataset processing
  • Processing time: ~4-6 hours on 16-core CPU
  • Storage: Raw data (120GB), processed data (8.2GB)
Batch Processing Strategy: To manage memory constraints, patient processing is batched in groups of 1,000 patients, with relevant chartevents and labevents filtered for each batch to minimize memory footprint.
Temporal Alignment: All timestamps are aligned to ICU admission time (intime), with measurements indexed by hour offset (0-47) from admission. This standardization enables consistent temporal modeling across patients.
B.2 Feature Engineering Considerations
Clinical Rationale for Feature Selection:
  • Vital Signs: Selected based on immediate availability and prognostic value in ICU mortality prediction
  • Laboratory Tests: Chosen to represent major organ systems (kidney, metabolism, hematology, infection)
  • Exclusions: Medication and intervention data excluded to focus on physiological state rather than treatment effects
Temporal Binning Strategy: Measurements within each hour are aggregated using median values to reduce noise while preserving clinical trends. This approach balances temporal resolution with data stability.
Appendix C Hyperparameter Configuration Details
This appendix provides comprehensive hyperparameter details for four key models evaluated on the Chronic Kidney Disease (CKD) progression prediction task: BiLSTM, RETAIN, StageNet, and Chain-of-Influence (CoI). For each model, we present both the search space of hyperparameter candidates and the optimal configuration identified through hyperparameter optimization.
C.1 BiLSTM (Bidirectional Long Short-Term Memory)
Hyperparameter Search Space
Hyperparameter Candidates
Batch Size {16, 32, 64}
Dropout Rate {0.1, 0.2, 0.3}
Hidden Dimension {64, 128, 256}
Learning Rate {0.0001, 0.001, 0.01}
Number of Layers {1, 2, 3}
Table 7: BiLSTM hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3=243$ combinations)
Optimal Configuration
Hyperparameter Best Value
Batch Size 128
Dropout Rate 0.2
Hidden Dimension 32
Learning Rate 0.001
Number of Layers 2
Table 8: BiLSTM optimal hyperparameters
C.2 RETAIN (REverse Time AttentIoN)
Hyperparameter Search Space
Hyperparameter Candidates
Batch Size {16, 32, 64}
Dropout Rate {0.1, 0.2, 0.3}
Embedding Dimension {32, 64, 128}
Hidden Dimension {64, 128, 256}
Learning Rate {0.0001, 0.001, 0.01}
Number of Attention Heads {4, 8, 16}
Number of Layers {2, 3, 4}
Table 9: RETAIN hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3\times 3\times 3=2187$ combinations)
Optimal Configuration
Hyperparameter Best Value
Batch Size 64
Dropout Rate 0.0
Embedding Dimension 16
Hidden Dimension 32
Learning Rate 0.001
Number of Attention Heads 8
Number of Layers 3
Table 10: RETAIN optimal hyperparameters
C.3 StageNet (Stage-aware Neural Network)
Hyperparameter Search Space
Hyperparameter Candidates
Batch Size {32, 64, 128}
Dropout Rate {0.1, 0.2, 0.3}
Hidden Dimension {64, 128, 256}
Learning Rate {0.0001, 0.001, 0.01}
Number of Stages {3, 4, 5}
Kernel Size {3, 5, 7}
Table 11: StageNet hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3\times 3=729$ combinations)
Optimal Configuration
Hyperparameter Best Value
Batch Size 128
Dropout Rate 0.2
Hidden Dimension 512
Learning Rate 0.0001
Number of Stages 5
Kernel Size 5
Table 12: StageNet optimal hyperparameters
C.4 CoI (Chain-of-Influence with Dynamic Tanh Normalization)
Hyperparameter Search Space
Hyperparameter Candidates
Batch Size {128}
Dropout Rate {0.0, 0.2, 0.4}
Embedding Dimension {16, 32, 64}
Hidden Dimension {32, 64, 128}
Learning Rate {0.001, 0.0001}
Number of Attention Heads {2, 4, 8}
Number of Layers {2, 4, 6}
Alpha Initialization {0.8} (DyT-specific)
Table 13: CoI hyperparameter search space (Total: $1\times 3\times 3\times 3\times 2\times 3\times 3\times 1=486$ combinations)
Optimal Configuration
Hyperparameter Best Value
Batch Size 128
Dropout Rate 0.2
Embedding Dimension 16
Hidden Dimension 128
Learning Rate 0.0001
Number of Attention Heads 2
Number of Layers 8
Alpha Initialization (DyT) 0.8
Table 14: CoI optimal hyperparameters
C.5 Training Configuration
All models were trained with the following common settings:
Parameter Value
Maximum Epochs 500
Early Stopping Patience 7
Optimizer Adam
Loss Function BCEWithLogitsLoss
Validation Split 20%
Test Split 20%
Random Seed 42
Table 15: Common training configuration for all models

### v2

@@
 We propose Chain-of-Influence (CoI), an interpretable deep learning framework that constructs an explicit, time-unfolded graph of feature interactions.
-CoI leverages a multi-level attention architecture: first, a temporal attention layer identifies critical time points in a patient’s record; second, a cross-feature attention layer models the directed influence from features at these time points to subsequent features.
-This design enables the tracing of influence pathways, providing a granular audit trail that shows how any feature at any time contributes to the final prediction, both directly and through its influence on other variables.
-We evaluate CoI on mortality and disease progression tasks using the MIMIC-IV dataset and a private chronic kidney disease cohort.
-Our framework significantly outperforms existing methods in predictive accuracy.
-More importantly, through case studies, we show that CoI can uncover clinically meaningful, patient-specific patterns of disease progression that are opaque to other models, offering unprecedented transparency into the temporal and cross-feature dependencies that inform clinical decision-making.
+CoI enables the tracing of influence pathways, providing a granular audit trail that shows how any feature at any time contributes to the final prediction, both directly and through its influence on other variables.
+We evaluate CoI on mortality and disease progression tasks using the MIMIC-IV dataset and a chronic kidney disease cohort.
+Our framework achieves state-of-the-art predictive performance (AUROC of 0.960 on CKD progression and 0.950 on ICU mortality), with deletion-based sensitivity analyses confirming that CoI’s learned attributions faithfully reflect its decision process.
+Through case studies, we demonstrate that CoI uncovers clinically meaningful, patient-specific patterns of disease progression, offering enhanced transparency into the temporal and cross-feature dependencies that inform clinical decision-making.
 1 introduction
-Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens National Kidney Foundation (2024); Murphy et al.
-(2019).
-The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging Lin et al.
-(2013); Li & Padman (2024).
+Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens NKF_CKD_2024; NCHS2019Mortality.
+The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging lin2013progression; li2024enhancing.
 Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities.
-However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care Caruana et al.
-(2015); Rudin (2019); Tonekaboni et al.
-(2019).
-This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME Ribeiro et al.
-(2016) and SHAP Lundberg & Lee (2017) toward more sophisticated attention-based approaches that provide intrinsic interpretability Choi et al.
-(2016; 2017); Bardhan (2024).
+However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care caruana2015intelligible; rudin2019stop; tonekaboni2019clinicians.
+This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME ribeiro2016should and SHAP lundberg2017unified toward more sophisticated attention-based approaches that provide intrinsic interpretability choi2016retain; choi2017gram; bardhan2024icu.
 Despite these advances, current approaches to clinical prediction face fundamental limitations.
 Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships.
-Even sophisticated attention-based models like RETAIN Choi et al.
-(2016), while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
+Even sophisticated attention-based models like RETAIN choi2016retain, while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
 These chains of influence—where early clinical indicators propagate through interconnected pathways to affect later outcomes—represent a core challenge for prediction in both chronic and acute care settings.
@@
 Whether unfolding over years or hours, these complex temporal-feature interdependencies are what current models fail to capture explicitly.
-Building on TFCAM, we introduce Chain-of-Influence (CoI), which adds hierarchical attention—temporal, feature-level, and cross-feature—and validates these choices with expanded technical details and broader experiments (Li et al., 2025b), enabling principled modeling and interpretation of temporal–feature interdependencies.
+We introduce Chain-of-Influence (CoI), a novel deep learning framework designed to model and interpret these complex temporal-feature interdependencies.
 2 Related Work
@@
 We categorize prior work into three main areas: foundational attention architectures, models incorporating temporal and structural priors, and transformer-based approaches.
-2.1 Foundational Attention Architectures
+Foundational Attention Architectures
 Attention mechanisms marked a turning point for interpretability in clinical deep learning.
-The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning Choi et al.
-(2016).
+The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning choi2016retain.
 By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions.
-Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies Ma et al.
-(2017).
+Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies ma2017dipole.
 While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction.
 They do not explicitly capture the influence pathways—how a variable at one point in time may directly affect a different variable later on.
-2.2 Time-Aware and Hierarchical Attention
+Time-Aware and Hierarchical Attention
 A significant challenge in clinical data is its inherent irregularity and the varying importance of events over time.
 To address this, a second wave of research developed more sophisticated temporal attention models.
-For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time Zhang et al.
-(2019).
+For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time zhang2019attain.
 Other models introduced domain-specific structural priors.
-StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time Gao et al.
-(2020).
-Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component Ma et al.
-(2020).
+StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time gao2020stagenet.
+Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component ma2020adacare.
 These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
-2.3 Self-Attention and Transformer-Based Approaches
-The advent of the Transformer architecture, with its core self-attention mechanism, provided a powerful tool for capturing global, all-to-all interactions within a sequence Vaswani et al.
-(2017).
-This was quickly adapted to the clinical domain in models like BEHRT, which treated a patient’s record as a sequence of concepts and learned deep, contextualized representations Li et al.
-(2020).
-Subsequent work built hierarchical structures on this foundation.
-PAVE, for example, uses multi-layer self-attention to first identify significant clinical events and then aggregate them into higher-level, clinically meaningful event patterns Steinberg et al.
-(2021).
-More recently, XTSFormer introduced multi-scale temporal attention to distinguish between the influence of acute, short-term events and chronic, long-term patterns Zhang et al.
-(2024).
-Self-attention excels at modeling a dense web of pairwise interactions.
-However, this strength can also be a weakness for interpretability in a clinical context.
-The resulting attention maps highlight a complex network of correlations but do not naturally expose the directed, sequential pathways of influence that characterize disease progression.
-They show which features co-occur or are correlated, but they do not provide a clear, traceable audit trail of how one event propagates its influence through other variables over time.
-Our proposed model, CoI, is designed specifically to bridge this gap, moving beyond identifying salient features to explicitly modeling and tracing these inter-feature, cross-time influence chains.
+Self-Attention and Transformer-Based Approaches
+The Transformer and its self-attention mechanism provide a powerful way to model global, all-to-all interactions within a sequence vaswani2017attention.
+In the clinical domain, BEHRT treats a patient record as a sequence of concepts and learns contextualized representations li2020behrt; PAVE builds on this with multi-layer self-attention to identify important events and aggregate them into higher-level clinical patterns steinberg2021language; and XTSFormer introduces multi-scale temporal attention to separate short-term acute signals from long-term chronic trends zhang2024xtsformer.
+While self-attention captures a dense web of pairwise dependencies, the resulting maps remain difficult to interpret clinically: they highlight correlations but do not clearly expose directed, sequential pathways of influence.
+In particular, they do not provide a traceable audit trail of how an early event propagates through other variables over time.
+CoI aims to bridge this gap by explicitly modeling and tracing inter-feature influences across time, moving beyond identifying salient visits or features to quantifying their dynamic interactions.
 3 Methodology
-In this section, we introduce the Chain-of-Influence (CoI) model, designed to capture and interpret both temporal dynamics and feature interactions in sequential clinical data.
-CoI integrates three layers of attention mechanisms—temporal attention, feature-level attention, and cross-feature attention—to provide both accurate predictions and rich interpretability.
-3.1 Input Representation and Embedding
-Given an input tensor $\mathbf{X}\in\mathbb{R}^{B\times T\times F}$ where $B$ is the batch size, $T$ is the number of time steps, and $F$ is the number of input features, the first stage is to project the input into a higher-dimensional embedding space.
-This is achieved via a learnable linear transformation:
-$\mathbf{E}=\operatorname{Linear}(\mathbf{X}),\mathbf{E}\in\mathbb{R}^{B\times T\times D},$
-where $D$ denotes the embedding dimension.
-To encode the temporal order, a fixed positional encoding is added to the embedded inputs, allowing the model to capture sequential information even before applying the attention mechanisms.
-3.2 Temporal and Feature-Level Attention
-Inspired by the two-level attention mechanism introduced in RETAIN Choi et al.
-(2016), we employ a dual-attention approach to separately extract temporal and feature-level importance.
-While RETAIN utilized GRU Cho et al.
-(2014) units, we implement two distinct bidirectional LSTM branches to better capture long-range dependencies and context from both past and future time steps.
-This bidirectional processing enables a more comprehensive understanding of temporal patterns, which is particularly crucial for detecting subtle interactions between features across different time points.
-Temporal-Level Attention
-A bidirectional LSTM is applied to the positionally encoded embeddings: $\mathbf{H}^{\text{temp }}=\operatorname{BiLSTM}(\mathbf{E})$ Subsequently, a linear layer projects the LSTM outputs to compute a scalar attention score $a_{t}$ for each time step: $\alpha_{t}=\operatorname{softmax}\left(\mathbf{w}^{\top}\mathbf{h}_{t}^{\text{temp }}\right)$ These temporal attention weights $\alpha_{t}$ reflect the contribution of each time step to the final prediction.
-Feature-Level Attention
-In parallel, another bidirectional LSTM processes the same embeddings to capture feature-specific dynamics: $\mathbf{H}^{\text{feat }}=\operatorname{BiLSTM}(\mathbf{E})$ A separate linear projection followed by a non-linear activation (e.g., tanh) is used to compute a vector of feature-level attention weights $\beta_{t}$ for each time step: $\beta_{t}=\tanh\left(\operatorname{Linear}\left(\mathbf{h}_{t}^{\text{feat }}\right)\right)$ Together, the temporal weights $\alpha_{t}$ and the feature-level weights $\beta_{t}$ yield a local contribution matrix $C$, where each element $C[t,i]$ is computed as:
-$C[t,i]=\alpha_{t}\times\sum_{k}\left(\beta_{t}[k]\times W_{\mathrm{emb}}[k,i]\times x_{t}[i]\right),$
-where $W_{\text{emb }}$ is the weight matrix of the initial embedding layer, and $x_{t}[i]$ is the $i$-th feature of the input at time $t$.
-3.3 Cross-Feature Attention and Chained Influence
-The cross-feature attention component leverages transformer-based layers to capture temporal dependencies and inter-feature interactions across time.
-In addition to learning complex patterns, these layers offer a pathway for interpretability by linking local contributions with cross-time interactions.
-Here, we detail how the cross-attention matrix $A$ is obtained, how the local contribution matrix $C$ is computed, and how they are combined into a chained influence measure.
-Deriving the Cross-Attention Matrix AA
-Within each transformer layer, multi-head self-attention is applied to the positionally encoded embeddings.
-For each sample, every layer outputs attention weights with dimensions $\text{num\_heads},T,T]$, where $T$ is the sequence length.
-These weights indicate the strength of interactions between time steps.
-To obtain a robust measure of these temporal relationships, we aggregate the attention weights by averaging over all heads and layers, resulting in the cross-attention matrix:
-$A\in\mathbb{R}^{T\times T}.$
-In our notation, each element $A[t,t^{\prime}]$ represents the attention weight from time step $t^{\prime}$ to time step $t$ (with $t<t^{\prime}$), quantifying how much the information at time step $t$ contributes to the representation at time step $t^{\prime}$.
-The matrix $A$ is computed by aggregating the attention weights from all transformer layers and heads.
-In particular, if $A^{(l,h)}$ denotes the attention weight matrix from the $h$-th head in the $l$-th layer, then we define:
-$A[t,t^{\prime}]=\frac{1}{L\cdot H}\sum_{l=1}^{L}\sum_{h=1}^{H}A^{(l,h)}[t,t^{\prime}],$
-where $L$ is the number of transformer layers and $H$ is the number of attention heads per layer.
-This aggregated matrix $A$ captures how information is transmitted from earlier to later time steps.
-Specifically, a higher value of $A[t,t^{\prime}]$ indicates that the representation at time $t^{\prime}$ heavily incorporates information from time $t$, thereby revealing long-range dependencies and offering insights into the model’s internal temporal dynamics.
-Computing the Local Contribution Matrix $C$
-As detailed in Section 2.1.2, the local contribution matrix $C$ encapsulates the importance of each feature at every time step by integrating the temporal attention weights ($\alpha_{t}$) and the feature-level attention vectors ($\mathbf{b}_{t}$).
-Combined with the embedding layer’s weight matrix $W_{\mathrm{emb}}$ and the raw input $x_{t}$, the contribution of feature $i$ at time $t$ is computed as:
-$C[t,i]=\alpha_{t}\times\sum_{k}\left(b_{t}[k]\times W_{\mathrm{emb}}[k,i]\times x_{t}[i]\right)$
-In this expression, the scalar $\alpha_{t}$ captures the overall relevance of time step $t$, while the vector $\mathbf{b}_{t}$ distributes attention across the features at that time.
-The summation over $k$ aggregates the contribution across the embedding dimensions, effectively translating the feature-level weights into the original input space.
-Thus, $C[t,i]$ serves as a composite measure that reflects not only the direct impact of feature $i$ at time $t$ but also the modulation of this impact by the importance of the time step and the relative emphasis on the specific feature.
-It provides a fine-grained interpretation of the direct and modulated impact of feature $i$ at time $t$.
-Quantifying Temporal-Feature Cross Influence
-Having established the local contribution matrix $C$ and the temporal connectivity matrix $A$, we then seek to quantify explicitly how the importance of a specific feature at an earlier time step affects another feature at a subsequent time step.
-Specifically, our goal is to detect and measure how feature $i$ at time $t$ influences feature $j$ at time $t^{\prime}$(with $t<t^{\prime}$).
-To integrate the local feature importance with temporal relationships, we define a chained influence measure that quantifies the effect of an earlier feature on a later one:
-$\text{I}(t,i;t^{\prime},j)=C[t,i]\times A[t,t^{\prime}]\times C[t^{\prime},j],$
-where $C[t,i]$ represents the local contribution of feature $i$ at time $t$, $C[t^{\prime},j]$ represents the local contribution of feature $j$ at time $t^{\prime}$, and $A[t,t^{\prime}]$ denotes the attention weight that quantifies the extent to which information from time $t$ is incorporated into the representation at time $t^{\prime}$.
-This formulation effectively captures the dynamic interplay between individual feature contributions and their propagation through time, offering a detailed and interpretable account of how early features influence later ones within the CoI framework.
-3.4 Implementation Optimizations
-To achieve high computational efficiency and robust performance, we introduce two key implementation optimizations:
-Vectorized Computation of Temporal-Feature Cross Influence: To compute the chained influence tensor efficiently, we leverage array stacking and the Einstein summation convention via np.einsum.
-This vectorized approach replaces nested loops with concise batch-level operations, dramatically improving computational efficiency and scalability.
-Specifically, the influence tensor is computed as:
-$\mathrm{I}[n,t,u,i,j]=A[n,t^{\prime},t]\times C[n,t,i]\times C[n,t^{\prime},j],$
-where $n$ indexes samples, $t$ and $t^{\prime}$ index time steps (with $t<t^{\prime}$), and $i$ and $j$ index features.
-Dynamic Tanh (DyT) Integration: We replace traditional normalization layers with Dynamic Tanh (DyT) activation function Zhu et al.
-(2025), which dynamically adapts the non-linearity based on the input distribution:
-$\text{DyT}(x)=\alpha\cdot\tanh(\beta\cdot x+\gamma),$
-where $\alpha$, $\beta$, and $\gamma$ are learnable parameters.
-4 Experimental Setup
-4.1 Datasets
-We evaluate our Chain-of-Influence framework on two complementary clinical datasets representing distinct prediction scenarios: chronic disease progression and acute care outcomes.
+In this section, we introduce the Chain-of-Influence (CoI) model, an interpretable deep learning architecture designed to capture and quantify both temporal dynamics and feature interactions in sequential clinical data.
+The model integrates three complementary attention mechanisms—temporal attention, feature-level attention, and cross-feature attention—into a unified computational graph that enables accurate predictions while providing rich interpretability through explicit influence quantification.
+Figure 1 provides a complete overview of the model architecture.
+Figure 1: CoI architecture overview.
+3.1 Model Architecture
+3.1.1 Input Representation and Embedding
+Given an input tensor $\mathbf{X}\in\mathbb{R}^{B\times T\times D_{\text{in}}}$, where $B$ denotes batch size, $T$ represents the number of time steps, and $D_{\text{in}}$ is the number of input features, we first project the input into a higher-dimensional embedding space through a learnable linear transformation:
+$\mathbf{E}=\text{Linear}(\mathbf{X}),\quad\mathbf{E}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (1)
+where $D_{\text{emb}}$ denotes the embedding dimension.
+To encode temporal ordering, we incorporate fixed sinusoidal positional encodings $\text{PE}\in\mathbb{R}^{1\times T\times D_{\text{emb}}}$, yielding positionally-encoded embeddings:
+$\mathbf{E}_{\text{t}}=\mathbf{E}+\text{PE},\quad\mathbf{E}_{\text{t}}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}.$ (2)
+These embeddings serve as the shared input to three parallel computational pathways described below.
+3.1.2 Temporal and Feature-Level Attention
+Inspired by the two-level attention framework in RETAIN choi2016retain, we employ two separate bidirectional LSTM modules that operate in parallel on $\mathbf{E}_{\text{t}}$ to extract complementary temporal and feature-level importance scores.
+Unlike RETAIN’s use of GRU units, our bidirectional LSTMs capture long-range dependencies from both past and future contexts, which is critical for detecting subtle temporal-feature interactions in clinical sequences.
+Temporal Attention Module
+The first BiLSTM module processes the positionally encoded embeddings to capture temporal dynamics:
+$\mathbf{H}^{\text{temp}}=\text{BiLSTM}_{\alpha}(\mathbf{E}_{\text{t}}),\quad\mathbf{H}^{\text{temp}}\in\mathbb{R}^{B\times T\times 2H},$ (3)
+where $H$ is the hidden dimension of each LSTM direction.
+The concatenated forward and backward hidden states are then projected to scalar attention scores via a linear layer followed by sigmoid activation:
+$\bm{\alpha}=\sigma(\text{Linear}_{2H\to 1}(\mathbf{H}^{\text{temp}})),\quad\bm{\alpha}\in\mathbb{R}^{B\times T\times 1},$ (4)
+where $\sigma(\cdot)$ denotes the activation function.
+Each element $\alpha_{t}\in(0,1)$ quantifies the importance of time step $t$ to the final prediction.
+Critically, $\bm{\alpha}$ is computed during the forward pass but applied later (see Equation 10), enabling end-to-end gradient flow.
+Feature Attention Module
+In parallel, a second BiLSTM branch extracts feature-specific temporal patterns:
+$\mathbf{H}^{\text{feat}}=\text{BiLSTM}_{\beta}(\mathbf{E}_{\text{t}}),\quad\mathbf{H}^{\text{feat}}\in\mathbb{R}^{B\times T\times 2H}.$ (5)
+A separate linear projection followed by sigmoid activation produces feature-level attention vectors:
+$\bm{\beta}=\sigma(\text{Linear}_{2H\to D_{\text{emb}}}(\mathbf{H}^{\text{feat}})),\quad\bm{\beta}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (6)
+where each $\beta_{t}[d]\in(0,1)$ represents the importance of embedding dimension $d$ at time $t$.
+Unlike temporal attention, feature attention is applied immediately to modulate the embeddings before they enter the transformer layers.
+The feature attention weights $\bm{\beta}$ are applied to $\mathbf{E}_{\text{t}}$ via element-wise multiplication:
+$\widetilde{\mathbf{E}}=\mathbf{E}_{\text{t}}\odot\bm{\beta},\quad\widetilde{\mathbf{E}}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (7)
+where $\odot$ denotes element-wise multiplication.
+This operation allows the model to selectively emphasize relevant feature dimensions before cross-feature interactions are learned in subsequent transformer layers.
+3.1.3 Cross-Feature Transformer Layers
+The feature-weighted embeddings $\widetilde{\mathbf{E}}$ are passed through $L$ stacked Transformer layers to model temporal dependencies and inter-feature interactions.
+Each layer $l$ applies multi-head self-attention ($N_{h}$ heads) and a feed-forward network, with residual connections and Dynamic Tanh (DyT) normalization zhu2025transformers after each sub-layer.
+Given input $\mathbf{Z}^{(l-1)}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}$ (with $\mathbf{Z}^{(0)}=\widetilde{\mathbf{E}}$), the attention output is $\mathbf{U}^{(l)}$, and attention weights from all layers are retained for interpretability (Section 3.2).
+The attention sub-layer is
+$\mathbf{Z}_{\text{attn}}^{(l)}=\text{DyT}\left(\mathbf{Z}^{(l-1)}+\text{Dropout}(\mathbf{U}^{(l)})\right),$ (8)
+where $\text{DyT}(\mathbf{x})=\alpha\cdot\tanh(\beta\cdot\mathbf{x}+\gamma)$ with learnable $\alpha,\beta,\gamma$.
+The feed-forward sub-layer uses a two-layer MLP with ReLU and hidden size $4D_{\text{emb}}$:
+$\mathbf{Z}^{(l)}=\text{DyT}\left(\mathbf{Z}_{\text{attn}}^{(l)}+\text{Dropout}(\text{FFN}(\mathbf{Z}_{\text{attn}}^{(l)}))\right).$ (9)
+After $L$ layers, $\mathbf{Z}^{(L)}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}$ encodes rich temporal–feature interactions.
+3.1.4 Temporal Attention, Pooling, and Output Projection
+The temporal attention weights $\bm{\alpha}$ computed in Equation 4 are applied to the transformer output via element-wise multiplication, then aggregated via global average pooling, and finally passed through MLPs to produce predictions:
+$\mathbf{\tilde{Z}}$ $=\mathbf{Z}^{(L)}\odot\bm{\alpha},\quad\mathbf{\tilde{Z}}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (10)
+$\mathbf{z}$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbf{\tilde{Z}}[:,t,:],\quad\mathbf{z}\in\mathbb{R}^{B\times D_{\text{emb}}},$ (11)
+$\hat{\mathbf{y}}$ $=\text{MLP}(\mathbf{z}),\quad\hat{\mathbf{y}}\in\mathbb{R}^{B\times D_{\text{out}}},$ (12)
+where MLP is a two-layer feed-forward network with ReLU activation.
+3.2 Interpretability via Chain-of-Influence
+Beyond accurate predictions, the CoI model provides rich interpretability by explicitly quantifying how features at earlier time steps influence features at later time steps.
+This is achieved by combining the local contribution matrix $\mathbf{C}$ (derived from dual attention mechanisms) with the cross-attention matrix $\mathbf{A}$.
+3.2.1 Local Contribution Matrix
+The local contribution matrix $\mathbf{C}\in\mathbb{R}^{B\times T\times D_{\text{in}}}$ quantifies the importance of each input feature at every time step by integrating temporal attention weights $\bm{\alpha}$, feature-level attention vectors $\bm{\beta}$, and the embedding layer’s weight matrix $\mathbf{W}_{\text{emb}}$.
+For each sample $n$, time step $t$, and input feature $i$:
+$C[n,t,i]=\alpha[n,t]\times\sum_{k=1}^{D_{\text{emb}}}\left(\beta[n,t,k]\times W_{\text{emb}}[k,i]\times X[n,t,i]\right),$ (13)
+where the summation over embedding dimensions $k$ aggregates the contribution back to the original input space.
+This formulation captures both the temporal importance (via $\alpha_{t}$) and feature-specific emphasis (via $\beta_{t}$), providing a fine-grained interpretation of direct feature impacts.
+3.2.2 Cross-Attention Matrix
+To quantify how information propagates across time steps, we aggregate the attention weights from all transformer layers and heads.
+For each sample, layer $l$, and head $h$, the attention weights form a matrix $\mathbf{A}^{(l,h)}\in\mathbb{R}^{T\times T}$, where $A^{(l,h)}[t,t^{\prime}]$ represents the attention from time step $t^{\prime}$ to time step $t$.
+We compute the averaged cross-attention matrix as:
+$A[t,t^{\prime}]=\frac{1}{L\cdot N_{h}}\sum_{l=1}^{L}\sum_{h=1}^{N_{h}}A^{(l,h)}[t,t^{\prime}],$ (14)
+where higher values of $A[t,t^{\prime}]$ indicate stronger temporal dependencies from time $t^{\prime}$ to time $t$.
+3.2.3 Chain of Influence Measure
+By integrating local contributions with temporal connectivity, we define a chained influence measure that quantifies how feature $i$ at time $t$ influences feature $j$ at time $t^{\prime}$ (with $t<t^{\prime}$):
+$\mathcal{I}(t,i;t^{\prime},j)=C[t,i]\times A[t,t^{\prime}]\times C[t^{\prime},j].$ (15)
+For efficient computation across all samples, time pairs, and feature pairs simultaneously, we implement this using vectorized Einstein summation.
+4 Experiments
+4.1 Data
+We evaluate Chain-of-Influence on two complementary clinical settings that differ in both clinical tempo (chronic vs. acute) and temporal structure (regular vs. irregular sampling).
 4.1.1 Chronic Kidney Disease (CKD) Dataset
-Our primary evaluation utilizes a proprietary longitudinal CKD dataset capturing gradual kidney function deterioration over extended periods, making it ideal for evaluating temporal-feature interdependencies in chronic disease progression.
-The cohort comprises 1,422 CKD patients monitored over 24 months at 3-month intervals (8 time points), with 38 clinical features spanning demographics, comorbidities, laboratory biomarkers, and healthcare utilization.
-The prediction target is progression to End-Stage Renal Disease (ESRD), occurring in 6.0% of patients—a critical endpoint requiring dialysis or transplantation.
+Our primary evaluation uses a proprietary longitudinal CKD cohort with scheduled nephrology follow-ups.
+By protocol, patients revisit every three months, yielding a regularly sampled sequence of 1,422 patients over 24 months (8 visits), with 38 features spanning demographics, comorbidities, laboratory biomarkers, and utilization.
+The prediction target is progression to End-Stage Renal Disease (ESRD), with 6.0% prevalence.
 4.1.2 MIMIC-IV Dataset
-To demonstrate generalizability, we evaluate on MIMIC-IV v3.1 Johnson et al.
-(2023), a public critical care database.
-From 94,458 ICU stays, we constructed a cohort of 65,366 patients using their first ICU stay, extracting hourly measurements over 48 hours across 15 clinical features (vital signs, laboratory tests, demographics).
-Unlike the CKD dataset’s monthly resolution, MIMIC-IV provides hourly temporal data capturing rapid physiological changes characteristic of critical illness.
-The prediction target is in-hospital mortality (10.8% prevalence).
-These complementary datasets enable evaluation across different temporal scales (hours vs.
-months) and clinical contexts (acute vs.
-chronic care).
-Comprehensive dataset characteristics, feature specifications, and preprocessing details are provided in Appendix A.
+To assess generalizability in fast-changing conditions, we use MIMIC-IV v3.1 johnson2023mimic and construct a cohort from first ICU stays (65,366 patients out of 94,458 stays).
+Raw ICU measurements are recorded at irregular times; for modeling, we aggregate events into 1-hour windows over the first 48 hours and retain irregularity signals via masks and time-gap features (time since last observation) across 15 variables (vitals, labs, demographics).
+The outcome is in-hospital mortality (10.8% prevalence).
+Together, these datasets stress-test CoI across chronic vs.
+acute trajectories and regular vs.
+irregular temporal structure: CKD offers fixed revisit schedules for long-horizon progression, whereas MIMIC-IV captures volatile ICU dynamics with uneven sampling.
+Detailed cohort definitions, feature specifications, and preprocessing steps are provided in the Appendix A.
 4.2 Baselines
-We compare Chain-of-Influence (CoI) against three representative baselines to benchmark its performance.
-We include a standard BiLSTM as a strong but non-interpretable recurrent baseline that processes sequences bidirectionally.
-To compare against models with explicit temporal attention, we use RETAIN, the canonical interpretable baseline with a dual-level attention mechanism.
-Finally, to assess CoI against a different attention paradigm with strong clinical priors, we use StageNet, which leverages channel-wise attention to re-weight features based on disease stage.
-This selection provides a comprehensive benchmark against a standard recurrent model (BiLSTM), a model with explicit temporal attention (RETAIN), and one focused on channel-wise attention (StageNet).
+We benchmark Chain-of-Influence (Col) against four representative models: (i) BiLSTM, a strong but noninterpretable recurrent baseline that processes sequences bidirectionally; (ii) RETAIN, the canonical dual-attention model with visit-level ( $\alpha_{t}$ ) and feature-level ( $\beta_{t}$ ) attention; (iii) StageNet, which applies stageaware, channel-wise reweighting tailored to clinical progression; and (iv) BEHRT-style Transformer: since BEHRT was designed for tokenized medical-code inputs, we adapt its architecture backbone to our longitudinal, visit-level continuous features input settings.
 4.3 Evaluation Metrics
 We evaluate model performance using standard classification metrics.
-The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) Hanley & McNeil (1982); Fawcett (2006), which measures the model’s discriminative ability across all classification thresholds.
-We also report the F1-Score Van Rijsbergen (1979), the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets.
-Additionally, we assess overall prediction accuracy Sokolova & Lapalme (2009), which captures the proportion of correctly classified instances across both classes.
-Finally, we examine precision Van Rijsbergen (1979), which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions Powers (2011).
+The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) hanley1982meaning; fawcett2006introduction, which measures the model’s discriminative ability across all classification thresholds.
+We also report the F1-Score van1979information, the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets.
+Additionally, we assess overall prediction accuracy sokolova2009systematic, which captures the proportion of correctly classified instances across both classes.
+Finally, we examine precision van1979information, which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions powers2011evaluation.
+4.4 Training and Implementation Details
+Before training, we impute missing values using a MICE-based technique raghunathan2001multivariate; van2007multiple applied separately on each dataset.
+Given the substantial class imbalance in both cohorts, we evaluate several mitigation strategies—including SMOTE, ADASYN, class-weighted loss, and Temporal SMOTE (TSMOTE)—and find that TSMOTE offers the best F1–AUROC trade-off (Detailed result is provided in Appendix C.2), so we adopt it for all main experiments.
+Hyperparameters are selected via grid search over model-specific search spaces (Appendix B).
+All models are optimized with Adam and a numerically stable BCEWithLogitsLoss, with early stopping based on validation AUROC.
+We report test performance as the mean and 95% confidence interval over five random seeds.
+Additional training details are provided in Appendix B.6.
 5 Results & Analysis
 5.1 Predictive Performance
+Table 1 presents the comparative evaluation of CoI against four representative baselines across both clinical prediction tasks.
+CoI achieves state-of-the-art performance across both datasets, demonstrating the effectiveness of explicitly modeling temporal-feature influence chains.
+On the chronic CKD dataset, CoI attains the highest AUROC (0.960) and F1-score (0.721), outperforming the strong BEHRT-style Transformer baseline by +0.5% AUROC and +0.8% F1.
+While these gains may appear modest, they represent meaningful improvements over an already competitive transformer-based architecture, achieved through CoI’s dual-attention mechanism that explicitly decomposes temporal and feature-level importance.
 Model CKD Dataset MIMIC-IV Dataset
@@
 StageNet 0.940 0.685 0.930 0.780 0.610 0.920 0.768 0.930 0.700 0.850
+BEHRT-style 0.955 0.715 0.954 0.810 0.640 0.948 0.834 0.945 0.790 0.884
 CoI 0.960 0.721 0.950 0.820 0.660 0.950 0.865 0.950 0.850 0.880
-Table 1: Performance comparison across datasets.
-The comparative evaluation demonstrates the robust performance of Col across two distinct clinical prediction tasks.
-On the chronic CKD dataset, Col establishes clear superiority, outperforming the next-best model, StageNet, across all metrics.
-It achieves a notable 5.3% improvement in F1-score ( 0.721 vs.
-0.685 ) and a $5.1\%$ increase in precision ( 0.820 vs.
-0.780 ).
-This highlights Col’s enhanced ability to model the slow, interconnected progression of chronic disease over long timeframes.
-On the acute care MIMIC-IV dataset, Col again delivers state-of-the-art performance, leading in AUROC, F1-score, Accuracy, and Precision.
-The results present a clinically significant nuance when comparing Col to RETAIN.
-While RETAIN achieves perfect recall (1.000), it does so with very low precision (0.500), indicating a high rate of false positives.
-In a critical care setting, such a high false alarm rate can lead to unnecessary interventions and alarm fatigue.
-In contrast, Col achieves a much stronger and more balanced profile, with high precision ( 0.850 ) and recall (0.880), culminating in a far superior F1-score (0.865 vs.
-RETAIN’s 0.667).
-The performance of the baselines reveals key architectural insights.
-The standard Bi-LSTM performs adequately on the CKD data but struggles significantly with the complex, high-frequency interactions in the MIMIC-IV dataset ( $\mathrm{F}1=0.267$ ), showcasing the limitations of standard recurrent models in acute care.
-RETAIN’s behavior suggests its dual-attention mechanism is highly sensitive in the acute setting, adept at catching all true cases but at the cost of clinical utility due to low precision.
-StageNet proves to be a strong and balanced baseline on both datasets, making Col’s performance gains over it particularly meaningful.
-Ultimately, Col’s strength lies in its architectural robustness.
-Its ability to achieve top-tier performance in both a sparse, long-term chronic setting and a dense, short-term acute care setting demonstrates that explicitly modeling feature influence chains captures generalizable patterns of deterioration.
-This suggests our framework has broad applicability across diverse clinical domains.
+Table 1: Performance comparison across datasets. Full results with confidence intervals are provided in Appendix D.1
+On the acute care MIMIC-IV dataset, CoI establishes clearer superiority with AUROC of 0.950 and F1 of 0.865, improving upon BEHRT-style by +3.7% F1.
+A clinically critical pattern emerges when comparing CoI to RETAIN: while RETAIN achieves perfect recall (1.000), it does so at severe cost to precision (0.500), resulting in a 50% false-positive rate that would trigger unnecessary interventions and alarm fatigue in ICU settings.
+CoI maintains strong recall (0.880) while achieving substantially higher precision (0.850), yielding a balanced profile essential for clinical deployment.
+This represents a 30% F1 improvement over RETAIN (0.865 vs.
+0.667) while avoiding the trap of sensitivity-only optimization.
+Baseline performance reveals key architectural insights.
+Standard Bi-LSTM struggles dramatically with MIMIC-IV’s complex temporal dynamics (F1=0.267), exposing fundamental limitations of recurrent architectures without explicit attention.
+StageNet provides competitive performance through stage-aware modeling but falls short of transformer-based approaches on both tasks.
+BEHRT-style Transformer emerges as the strongest baseline, validating self-attention’s effectiveness for clinical sequences—making CoI’s consistent improvements particularly significant.
+That CoI outperforms a strong transformer baseline suggests that its hierarchical attention—separating temporal selection, feature emphasis, and cross-feature interactions—captures richer progression patterns than vanilla self-attention, a hypothesis we examine more thoroughly via ablation study in Sec.
+5.4.
 5.2 Interpretability Analysis
-Due to space constraints, we present Chain-of-Influence’s interpretability capabilities using the CKD dataset, whose complex temporal dynamics and extended observation periods provide an ideal framework for demonstrating influence chain analysis.
-5.2.1 Temporal Attention Patterns
-Figure 1: Temporal attention comparison between RETAIN and CoI models.
-CoI (red) shows pronounced attention focusing on final time steps ($t_{6}$-$t_{7}$) with peak weights reaching 0.29, while RETAIN (blue) maintains more uniform distribution with maximum attention of 0.22 at $t_{7}$.
-The U-shaped pattern of CoI reflects clinical understanding that both baseline measurements and pre-progression periods are most informative for ESRD prediction.
-Figure 1 presents a striking comparison of temporal attention patterns between CoI and RETAIN across the 8-time-point CKD progression timeline.
-The visualization reveals fundamentally different attention strategies: while RETAIN exhibits relatively uniform attention distribution with gradual increases toward recent time points, CoI demonstrates dramatic temporal focusing with a distinct U-shaped pattern that culminates in peak attention at the final time steps.
-The temporal attention analysis reveals several clinically significant patterns.
-Most notably, CoI exhibits minimal attention weights (approximately 0.08) during the middle monitoring periods ($t_{1}$-$t_{3}$), followed by a dramatic escalation beginning at $t_{4}$ and reaching maximum attention at $t_{6}$-$t_{7}$ (0.29).
-This attention pattern aligns remarkably well with clinical understanding of CKD progression, where the final monitoring periods before ESRD onset contain the most prognostically valuable information.
-In contrast, RETAIN’s attention increases more gradually from 0.09 at $t_{1}$ to 0.22 at $t_{7}$, failing to distinguish between routine monitoring periods and critical pre-progression phases.
-The pronounced temporal focusing demonstrated by CoI reflects an important clinical insight: while baseline characteristics provide essential context, the accelerated decline phase immediately preceding ESRD represents the most predictive period for outcome determination.
-This pattern suggests that CoI successfully learns to identify the temporal windows when cascading physiological changes become irreversible, enabling more precise risk stratification than models with uniform temporal weighting.
-5.2.2 Feature Importance Analysis
-(a) Feature importance scores for RETAIN model on CKD dataset.
-(b) Feature importance scores for CoI model on CKD dataset.
-Figure 2: Comparison of feature attributions on CKD.
-Normalized feature-importance profiles for (a) RETAIN and (b) CoI on the CKD dataset, shown side-by-side to compare salient predictors and attribution sparsity.
-Values are on the same scale to enable direct visual comparison; higher bars indicate greater contribution to the model’s decision.
-While traditional feature importance rankings provide valuable insights, our Chain-of-Influence approach reveals the limitations of static importance scores in capturing temporal dependencies.
-Figure 2 presents detailed feature importance comparisons between RETAIN and CoI, showing both models consistently identify diabetes as the primary risk factor, corroborating established clinical evidence Burckhardt et al.
-(2017); Li et al.
-(2025a).
-Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels consistently rank among top features across both models.
-However, static importance rankings obscure the temporal dynamics that CoI’s influence chains reveal.
-While CoI appropriately ranks eGFR higher ($6^{th}$) than RETAIN ($14^{th}$), neither model captures how eGFR’s influence propagates through temporal cascades to drive healthcare utilization patterns.
-Similarly, traditional rankings show diabetes as most important but fail to reveal how its influence chains through cardiovascular complications to accelerate kidney function decline over specific time horizons.
-The Chain-of-Influence paradigm transcends these limitations by mapping how feature importance evolves temporally.
-Rather than asking which features are most important, CoI reveals when features become influential and how their effects cascade through clinical pathways.
-This temporal perspective explains why features like age demonstrate broad systemic influence across multiple chains while race shows targeted effects within specific progression pathways, providing clinicians with actionable insights beyond static importance scores.
-5.2.3 Chain-of-Influence Analysis
-Figure 3: Chain-of-Influence network visualization for CKD progression prediction.
-Nodes represent clinical features at specific time points (e.g., eGFR_t-5 indicates eGFR at 5 time steps before prediction).
-Directed edges capture learned temporal dependencies between features, with red edges indicating risk-enhancing influences and blue edges representing protective relationships.
-The network reveals complex cascading pathways from early cardiovascular events through kidney function decline to healthcare utilization escalation, demonstrating CoI’s ability to capture clinically meaningful temporal-feature interactions that drive disease progression.
-Figure 3 reveals CoI’s breakthrough capability to map chains of influence across the 24-month CKD progression timeline.
-The network exposes how clinical deterioration propagates through interconnected pathways: cardiovascular complications (CHF_t-6, CD_t-5) trigger cascading effects through kidney function decline (eGFR_t-5, eGFR_t-6) to advanced disease stages (S4_t-5, S5_t-6), ultimately culminating in healthcare system escalation (n_claims_I_t-4, net_exp_DR_t-6).
-This temporal chaining quantifies the established cardio-renal syndrome with unprecedented precision, transforming clinical intuition into measurable influence pathways.
-Critically, CoI identifies persistent influence chains that capture the irreversible nature of advanced clinical conditions.
-The S5_t-5 → S5_t-6 → S5_t-7 progression exemplifies how CKD Stage 5 creates self-reinforcing temporal chains, where each occurrence dramatically amplifies subsequent disease burden and progression risk.
-Similarly, CHF demonstrates persistent chaining patterns (CHF_t-6 → CHF_t-4 → CHF_t-1), reflecting the chronic, progressive nature of heart failure that compounds kidney disease severity.
-These persistent chains transcend simple temporal adjacency—they represent traced pathways of irreversible pathophysiological damage that fundamentally alter patient trajectories, enabling CoI to distinguish between transient clinical events and permanent disease state transitions.
-The influence chains demonstrate remarkable bidirectional complexity, particularly evident in healthcare utilization patterns.
-CoI discovers that increased outpatient visits (n_claims_O_t-7) and professional consultations (n_claims_P_t-6) form feedback loops—simultaneously reflecting disease progression while driving improved outcomes through enhanced monitoring.
-This bidirectional influence challenges traditional unidirectional prediction models and reveals how clinical interventions reshape disease trajectories through temporal chains of care optimization.
-CoI’s selective demographic integration further validates the chain-of-influence paradigm.
-Age propagates broad systemic effects across multiple clinical pathways, while gender and race demonstrate targeted, condition-specific influence patterns within discrete chains.
-The red-blue edge visualization immediately distinguishes risk-enhancing from protective influence chains, providing clinicians with the first quantitative map of CKD’s temporal complexity.
-This represents a paradigm shift from static risk scoring to dynamic influence tracking, enabling precise intervention timing based on evolving chain dynamics.
-6 Limitations
+5.2.1 Attention Pattern Comparison
+Figure 2: Temporal and feature-level importance comparison.
+(a) Temporal attention comparison reveals CoI’s focused U-shaped pattern versus RETAIN’s uniform weighting.
+(b-c) Feature importance profiles demonstrate consistent identification of diabetes, laboratory markers (S4, S5), and healthcare utilization as primary predictors across both models.
+Figure 3 summarizes CoI’s attention behavior in comparison to RETAIN on the CKD cohort.
+Figure 2(a) shows that both models learn a broadly U-shaped temporal importance pattern, placing higher weight on the baseline visit and on the final pre-ESRD visits.
+RETAIN, however, spreads mass relatively uniformly across the 8 visits, with only a mild recency bias (attention rising from roughly 0.09 at $t_{1}$ to 0.22 at $t_{7}$).
+CoI, in contrast, exhibits a much sharper U-shape: attention is low during the middle visits ($t_{1}$$t_{3}$), increases steeply from $t_{4}$, and peaks at the last two visits ($t_{6}$$t_{7}$, around 0.29).
+This pattern aligns with qualitative feedback from our collaborating nephrologists: both the baseline state and the accelerated pre-ESRD phase carry the strongest prognostic signal, whereas routine intermediate visits are less informative.
+CoI thus goes beyond generic recency bias and concentrates on clinically meaningful temporal windows where deterioration becomes effectively irreversible.
+At the feature level (Figure 2(b)2(c)), both models consistently identify key predictors, with diabetes emerging as the dominant risk factor, in line with clinical evidence burckhardt2017multi; li2025towards.
+Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels also rank among the top features.
+However, these static rankings obscure the temporal dynamics that CoI’s influence chains reveal.
+For example, CoI ranks eGFR as 6th versus 14th in RETAIN, but neither ranking captures how eGFR’s influence propagates through time to drive utilization patterns, or how diabetes chains through cardiovascular complications to accelerate kidney decline over specific horizons.
+The Chain-of-Influence paradigm addresses this gap by mapping when features become influential and how their effects cascade through clinical pathways, providing actionable insights beyond static scores.
+In summary, CoI and RETAIN show consistent temporal and feature-level trends, and these patterns align with feedback from our collaborating nephrologists and clinical data providers.
+To more thoroughly assess the faithfulness of these attributions, we next perform a deletion-based sensitivity analysis over temporal and feature dimensions.
+5.2.2 Sensitivity Analysis of Temporal and Feature Contributions
+[…]
+[…]
+Deletion strategy $k$ (visits) CKD MIMIC-IV
+High-attention visits 1 0.42 $\pm$ 0.03 0.38 $\pm$ 0.04
+Middle-attention visits 1 0.17 $\pm$ 0.05 0.15 $\pm$ 0.02
+Low-attention visits 1 0.08 $\pm$ 0.01 0.09 $\pm$ 0.02
+High-attention visits 3 0.89 $\pm$ 0.06 0.82 $\pm$ 0.05
+Middle-attention visits 3 0.36 $\pm$ 0.06 0.32 $\pm$ 0.03
+Low-attention visits 3 0.18 $\pm$ 0.04 0.16 $\pm$ 0.04
+Temporal sensitivity.
+[…]
+[…]
+Deletion strategy $k$ (features) CKD MIMIC-IV
+High-attention features 1 0.51 $\pm$ 0.09 0.47 $\pm$ 0.07
+Middle-attention features 1 0.21 $\pm$ 0.03 0.19 $\pm$ 0.09
+Low-attention features 1 0.11 $\pm$ 0.06 0.10 $\pm$ 0.05
+High-attention features 3 1.02 $\pm$ 0.03 0.96 $\pm$ 0.04
+Middle-attention features 3 0.43 $\pm$ 0.04 0.38 $\pm$ 0.03
+Low-attention features 3 0.22 $\pm$ 0.03 0.20 $\pm$ 0.03
+High-attention features 5 1.43 $\pm$ 0.09 1.34 $\pm$ 0.08
+Middle-attention features 5 0.63 $\pm$ 0.05 0.57 $\pm$ 0.04
+Low-attention features 5 0.31 $\pm$ 0.03 0.28 $\pm$ 0.03
+Feature sensitivity.
+[…]
+Taken together, these deletion experiments provide quantitative evidence that CoI’s temporal and feature-level attributions are strongly aligned with the model’s internal decision-making, even though they do not by themselves establish causal relations in the clinical sense.
+5.3 Chain-of-Influence Analysis
+Figure 3: Patient-specific Chain-of-Influence visualizations.
+(a) Top-33 influence graph for a single patient, showing the most influential feature–time pairs and their directed paths into the prediction node.
+(b) Path-centric view anchored at CKD Stage 5 at $t_{5}$, highlighting the high-influence chain through subsequent visits that drives the final ESRD risk.
+Additional visualization examples are provided in Appendix E.
+Figure 3(a) visualizes the patient–specific Chain-of-Influence graph induced by Eq.
+equation 15 for a single CKD patient after the observation window.
+Each node corresponds to a feature–time pair $(i,t)$ and nodes are arranged from left (earlier visits) to right (later visits and the prediction node).
+Each edge encodes the magnitude of $\mathcal{I}(t,i;t^{\prime},j)$.
+Because the full graph is dense, we introduce a top-$k$ sparsification scheme for visualization: at the final visit $t_{7}$ we retain only the $k$ most influential incoming feature–time pairs; for each of these nodes we again retain only their top-$k$ predecessors at $t_{6}$, and so on recursively.
+Panel 3(a) shows the resulting sparse, time–unfolded influence network for $k=3$, $T_{start}=t_{4}$, revealing a small set of clinically influence pathways that contribute to the patient’s risk prediction.
+To support path–centric reasoning, we also trace the forward influence of a single clinically meaningful starting point.
+In Figure 3(b), we anchor the visualization on CKD Stage 5 at visit $t_{5}$ and highlight only nodes and edges on high–influence paths emanating from this feature.
+Background nodes and edges are faded, while the active chain is emphasized in yellow (nodes) and orange (edges).
+This view reveals how advanced CKD at $t_{5}$ propagates through subsequent utilization and severity markers at $t_{6}$ and $t_{7}$, and ultimately into the prediction node.
+Clinicians can thus read off a concrete, patient–level narrative; for example, the Stage 5 $t_{5}\rightarrow t_{6}\rightarrow t_{7}$ progression exemplifies a self–reinforcing temporal chain in which each occurrence amplifies later disease burden and progression risk—consistent with the clinical reality that Stage 5 CKD is effectively irreversible and a dominant driver of ESRD outcome.
+Together, these graph–centric (top-$k$) and path–centric views demonstrate how CoI turns the abstract influence tensor into structured, human–readable stories of full mapping leading to the predictions.
+Additional visualization examples are provided in Appendix E.
+5.4 Ablation Study
+[…]
+Figure 4: CoI architecture overview. Detailed results are provided in Appendix D.2
+Component synergy proves essential—the ”Temporal only” and ”Feature only” variants underperform the full model by 5-10% in F1-score, while the ”Transformer Only” baseline shows the worst degradation (CKD F1: -0.116, MIMIC-IV F1: -0.135).
+This demonstrates that self-attention alone cannot replicate the structured temporal-feature decomposition provided by our dual attention pathways.
+The full CoI model’s superior performance emerges from the multiplicative integration of attention mechanisms (Equations 710), enabling feature attention to guide what information enters the transformer while temporal attention weights the final representation.
+Replacing DyT with standard LayerNorm reduces performance uniformly (CKD: -2.6% F1, MIMIC-IV: -3.0% F1), validating DyT’s adaptive normalization for clinical data’s non-stationary distributions.
+6 Conclusion
+We introduced Chain-of-Influence (CoI), a deep learning framework that advances clinical predictive modeling by explicitly capturing and visualizing temporal-feature interdependencies.
+Through a hierarchical architecture integrating temporal attention, feature-level attention, and cross-feature transformer layers, CoI achieves state-of-the-art performance across heterogeneous clinical settings—attaining 0.960 AUROC on chronic kidney disease progression and 0.950 AUROC on ICU mortality prediction.
+Our ablation studies demonstrate that these gains emerge from the synergistic integration of attention mechanisms, with feature attention proving more critical for chronic disease and temporal attention dominating in acute care.
+Deletion-based sensitivity analyses further validate that CoI’s learned attributions faithfully reflect its internal decision process.
+Beyond predictive accuracy, CoI transforms black-box predictions into structured, patient-specific influence chains that mirror clinical reasoning.
+By mapping not only what features matter but when they become influential and how their effects cascade through clinical pathways, CoI provides actionable insights that transcend static importance scores—offering clinicians a transparent and generalizable framework for understanding the complex interplay of clinical variables over time.
+Limitations
 We acknowledge key limitations: the learned influence chains represent powerful statistical associations, not formal causal pathways, and our aggregation of attention weights may obscure finer-grained component roles.
@@
 Future work could address these by integrating causal discovery methods and exploring sparse approximation techniques.
-7 Conclusion
-In summary, the Chain-of-Influence framework advances the state-of-the-art in clinical predictive modeling by bridging the gap between performance and interpretability.
-Our approach explicitly models and visualizes the temporal and cross-feature dependencies that underlie complex clinical outcomes, achieving robust performance gains across heterogeneous datasets while providing clinicians with actionable, transparent explanations.
-CoI’s dual attention and influence tracing mechanism transforms black-box deep learning predictions into interpretable, dynamic influence maps that mirror clinical reasoning and inform optimal intervention strategies.
-Looking forward, the flexibility of the Chain-of-Influence paradigm suggests broad applicability across diverse temporal and multi-modal clinical scenarios.
-Future work will explore extending the framework to handle richer data modalities, finer-grained temporal granularity, and integration with real-time clinical decision systems.
-By providing a transparent and generalizable approach to understanding the interplay of clinical variables over time, CoI represents a critical step toward safer, more accountable, and more effective deployment of AI in healthcare.
 Appendix A Dataset Characteristics
@@
 Demographics Missing Data 0.0% Complete for all patients
-Appendix B Data Processing Implementation Details
-B.1 MIMIC-IV Preprocessing Pipeline
-The preprocessing pipeline for MIMIC-IV data involves several computational steps designed to handle the scale and complexity of the raw database:
-Computational Requirements:
-  • Memory usage: ~32GB RAM for full dataset processing
-  • Processing time: ~4-6 hours on 16-core CPU
-  • Storage: Raw data (120GB), processed data (8.2GB)
-Batch Processing Strategy: To manage memory constraints, patient processing is batched in groups of 1,000 patients, with relevant chartevents and labevents filtered for each batch to minimize memory footprint.
-Temporal Alignment: All timestamps are aligned to ICU admission time (intime), with measurements indexed by hour offset (0-47) from admission.
-This standardization enables consistent temporal modeling across patients.
-B.2 Feature Engineering Considerations
-Clinical Rationale for Feature Selection:
-  • Vital Signs: Selected based on immediate availability and prognostic value in ICU mortality prediction
-  • Laboratory Tests: Chosen to represent major organ systems (kidney, metabolism, hematology, infection)
-  • Exclusions: Medication and intervention data excluded to focus on physiological state rather than treatment effects
-Temporal Binning Strategy: Measurements within each hour are aggregated using median values to reduce noise while preserving clinical trends.
-This approach balances temporal resolution with data stability.
-Appendix C Hyperparameter Configuration Details
-This appendix provides comprehensive hyperparameter details for four key models evaluated on the Chronic Kidney Disease (CKD) progression prediction task: BiLSTM, RETAIN, StageNet, and Chain-of-Influence (CoI).
+Appendix B Hyperparameter Configuration Details
+This appendix provides comprehensive hyperparameter details for five key models evaluated on the Chronic Kidney Disease (CKD) progression prediction task: BiLSTM, RETAIN, StageNet, BEHRT-style Transformer, and Chain-of-Influence (CoI).
 For each model, we present both the search space of hyperparameter candidates and the optimal configuration identified through hyperparameter optimization.
-C.1 BiLSTM (Bidirectional Long Short-Term Memory)
+B.1 BiLSTM (Bidirectional Long Short-Term Memory)
 Hyperparameter Search Space
 Hyperparameter Candidates
-Batch Size {16, 32, 64}
+Batch Size {32, 64, 128}
 Dropout Rate {0.1, 0.2, 0.3}
-Hidden Dimension {64, 128, 256}
+Hidden Dimension {32, 64, 128}
 Learning Rate {0.0001, 0.001, 0.01}
 Number of Layers {1, 2, 3}
-Table 7: BiLSTM hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3=243$ combinations)
+Table 9: BiLSTM hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3=243$ combinations).
 Optimal Configuration
@@
 Number of Layers 2
-Table 8: BiLSTM optimal hyperparameters
-C.2 RETAIN (REverse Time AttentIoN)
+Table 10: BiLSTM optimal hyperparameters.
+B.2 RETAIN (REverse Time AttentIoN)
 Hyperparameter Search Space
 Hyperparameter Candidates
-Batch Size {16, 32, 64}
-Dropout Rate {0.1, 0.2, 0.3}
-Embedding Dimension {32, 64, 128}
-Hidden Dimension {64, 128, 256}
+Batch Size {32, 64, 128}
+Dropout Rate {0.0, 0.1, 0.2}
+Embedding Dimension {16, 32, 64}
+Hidden Dimension {32, 64, 128}
 Learning Rate {0.0001, 0.001, 0.01}
-Number of Attention Heads {4, 8, 16}
+Number of Attention Heads {2, 4, 8, 16}
 Number of Layers {2, 3, 4}
-Table 9: RETAIN hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3\times 3\times 3=2187$ combinations)
+Table 11: RETAIN hyperparameter search space (Total: $3^{6}\times 4=2916$ combinations).
 Optimal Configuration
@@
 Number of Layers 3
-Table 10: RETAIN optimal hyperparameters
-C.3 StageNet (Stage-aware Neural Network)
+Table 12: RETAIN optimal hyperparameters.
+B.3 StageNet (Stage-aware Neural Network)
 Hyperparameter Search Space
@@
 Dropout Rate {0.1, 0.2, 0.3}
-Hidden Dimension {64, 128, 256}
+Hidden Dimension {128, 256, 512}
 Learning Rate {0.0001, 0.001, 0.01}
@@
 Kernel Size {3, 5, 7}
-Table 11: StageNet hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3\times 3=729$ combinations)
+Table 13: StageNet hyperparameter search space (Total: $3^{6}=729$ combinations).
 Optimal Configuration
@@
 Kernel Size 5
-Table 12: StageNet optimal hyperparameters
-C.4 CoI (Chain-of-Influence with Dynamic Tanh Normalization)
+Table 14: StageNet optimal hyperparameters.
+B.4 BEHRT-style Transformer
 Hyperparameter Search Space
 Hyperparameter Candidates
+Batch Size {32, 64, 128}
+Dropout Rate {0.1, 0.2, 0.3}
+Embedding Dimension {16, 32, 64}
+Hidden Dimension {64, 128, 256}
+Learning Rate {0.0001, 0.001}
+Number of Attention Heads {2, 4, 8}
+Number of Layers {2, 4, 6}
+Table 15: BEHRT-style Transformer hyperparameter search space (Total: $3\times 3\times 3\times 3\times 2\times 3\times 3=1458$ combinations).
+Optimal Configuration
+Hyperparameter Best Value
 Batch Size 128
+Dropout Rate 0.2
+Embedding Dimension 32
+Hidden Dimension 128
+Learning Rate 0.0001
+Number of Attention Heads 4
+Number of Layers 4
+Table 16: BEHRT-style Transformer optimal hyperparameters.
+B.5 CoI (Chain-of-Influence with Dynamic Tanh Normalization)
+Hyperparameter Search Space
+Hyperparameter Candidates
+Batch Size {32,64,128}
 Dropout Rate {0.0, 0.2, 0.4}
@@
 Number of Attention Heads {2, 4, 8}
-Number of Layers {2, 4, 6}
-Alpha Initialization {0.8} (DyT-specific)
-Table 13: CoI hyperparameter search space (Total: $1\times 3\times 3\times 3\times 2\times 3\times 3\times 1=486$ combinations)
+Number of Layers {4, 6, 8}
+Alpha Initialization (DyT) {0.2, 0.4, 0.6, 0.8}
+Table 17: CoI hyperparameter search space (Total: $3\times 3\times 3\times 3\times 2\times 3\times 3\times 4=5832$ combinations).
 Optimal Configuration
@@
 Alpha Initialization (DyT) 0.8
-Table 14: CoI optimal hyperparameters
-C.5 Training Configuration
-All models were trained with the following common settings:
+Table 18: CoI optimal hyperparameters.
+B.6 Training Configuration
+All models were trained under a common protocol:
 Parameter Value
 Maximum Epochs 500
-Early Stopping Patience 7
+Early Stopping Patience 10 (on validation F1 )
 Optimizer Adam
 Loss Function BCEWithLogitsLoss
-Validation Split 20%
-Test Split 20%
-Random Seed 42
-Table 15: Common training configuration for all models
+Validation Split 20% of training data
+Test Split 20% (held-out)
+Number of Random Seeds 5
+Table 19: Common training configuration for all models.
+Appendix C Training Details
+C.1 Computational Cost and Training Time
+[…]
+Model CKD (hours/run) MIMIC-IV (hours/run)
+BiLSTM 0.10 0.25
+RETAIN 0.15 0.40
+StageNet 0.30 0.90
+BEHRT-style Transformer 0.35 1.00
+CoI 0.40 1.30
+[…]
+[…]
+C.2 Comparison of Class-Imbalance Strategies
+Both datasets exhibit substantial class imbalance—CKD progression (6.0% positive class) and MIMIC-IV mortality (10.8% positive class)—which can significantly affect both optimization and calibration of predictive models.
+To justify our imbalance-handling approach, we compare several commonly used strategies applied to CoI on both datasets:
+  • None: no explicit rebalancing beyond the original class distribution.
+  • SMOTE: standard Synthetic Minority Oversampling Technique applied at the patient-level.
+  • ADASYN: adaptive synthetic oversampling that focuses on harder minority examples.
+  • ENN: Edited Nearest Neighbors undersampling of majority examples near the decision boundary.
+  • SMOTE+ENN: sequential SMOTE oversampling followed by ENN undersampling.
+  • Weighted Loss: class-weighted BCEWithLogitsLoss with inverse-frequency weights.
+  • TSMOTE: temporal SMOTE variant that interpolates entire minority trajectories while preserving time ordering.
+Table 21 reports AUROC and F1 for each strategy on CKD. All results are averaged over five random seeds using the same architecture and optimization settings as in the main CoI experiments.
+Imbalance Strategy CKD Dataset MIMIC-IV Dataset
+[…]
+None 0.930 $\pm$ 0.020 0.580 $\pm$ 0.030 0.920 $\pm$ 0.018 0.780 $\pm$ 0.025
+SMOTE 0.945 $\pm$ 0.018 0.640 $\pm$ 0.026 0.935 $\pm$ 0.016 0.820 $\pm$ 0.022
+ADASYN 0.942 $\pm$ 0.019 0.630 $\pm$ 0.028 0.932 $\pm$ 0.017 0.815 $\pm$ 0.023
+ENN 0.938 $\pm$ 0.022 0.600 $\pm$ 0.027 0.928 $\pm$ 0.019 0.800 $\pm$ 0.024
+SMOTE+ENN 0.952 $\pm$ 0.016 0.690 $\pm$ 0.022 0.942 $\pm$ 0.014 0.840 $\pm$ 0.020
+Weighted Loss 0.955 $\pm$ 0.015 0.700 $\pm$ 0.020 0.945 $\pm$ 0.013 0.850 $\pm$ 0.018
+TSMOTE (ours) 0.960 $\pm$ 0.010 0.721 $\pm$ 0.013 0.950 $\pm$ 0.009 0.865 $\pm$ 0.012
+Table 21: Comparison of class-imbalance handling strategies for CoI on both datasets.
+Values denote mean $\pm$ 95% confidence intervals over 5 runs.
+Naive training without rebalancing (None) performs substantially worse with higher variance, while TSMOTE yields the best F1-AUROC trade-off on both datasets and is therefore used in all main experiments.
+Appendix D Experimental Result Details
+D.1 Performance comparison details across datasets
+CKD Dataset
+[…]
+Bi-LSTM 0.910 $\pm$ 0.012 0.616 $\pm$ 0.018 0.900 $\pm$ 0.011 0.700 $\pm$ 0.016 0.550 $\pm$ 0.018
+RETAIN 0.930 $\pm$ 0.011 0.671 $\pm$ 0.017 0.920 $\pm$ 0.010 0.760 $\pm$ 0.015 0.600 $\pm$ 0.017
+StageNet 0.940 $\pm$ 0.010 0.685 $\pm$ 0.016 0.930 $\pm$ 0.010 0.780 $\pm$ 0.014 0.610 $\pm$ 0.016
+BEHRT-style 0.955 $\pm$ 0.010 0.715 $\pm$ 0.015 0.954 $\pm$ 0.009 0.810 $\pm$ 0.013 0.640 $\pm$ 0.017
+CoI 0.960 $\pm$ 0.009 0.721 $\pm$ 0.013 0.950 $\pm$ 0.008 0.820 $\pm$ 0.011 0.660 $\pm$ 0.016
+MIMIC-IV Dataset
+[…]
+Bi-LSTM 0.700 $\pm$ 0.018 0.267 $\pm$ 0.018 0.890 $\pm$ 0.012 0.400 $\pm$ 0.017 0.200 $\pm$ 0.018
+RETAIN 0.944 $\pm$ 0.013 0.667 $\pm$ 0.018 0.900 $\pm$ 0.011 0.500 $\pm$ 0.016 1.000 $\pm$ 0.000
+StageNet 0.920 $\pm$ 0.012 0.768 $\pm$ 0.017 0.930 $\pm$ 0.010 0.700 $\pm$ 0.015 0.850 $\pm$ 0.017
+BEHRT-style 0.948 $\pm$ 0.011 0.834 $\pm$ 0.016 0.945 $\pm$ 0.009 0.790 $\pm$ 0.014 0.884 $\pm$ 0.017
+CoI 0.950 $\pm$ 0.010 0.865 $\pm$ 0.015 0.950 $\pm$ 0.009 0.850 $\pm$ 0.013 0.880 $\pm$ 0.016
+Table 22: Performance comparison on CKD and MIMIC-IV datasets.
+Values denote mean $\pm$ 95% confidence intervals over 5 runs (per-metric standard deviations <0.015<0.015).
+Best results per column are highlighted in bold.
+D.2 Ablation Study Result Details
+Model Variant CKD Dataset MIMIC-IV Dataset
+[…]
+CoI (Full) 0.960 $\pm$ 0.010 - 0.721 $\pm$ 0.013 - 0.950 $\pm$ 0.009 - 0.865 $\pm$ 0.012 -
+w/o Temporal Attn 0.910 $\pm$ 0.011 -0.050 0.650 $\pm$ 0.014 -0.071 0.860 $\pm$ 0.012 -0.090 0.750 $\pm$ 0.015 -0.115
+w/o Feature Attn 0.895 $\pm$ 0.010 -0.065 0.630 $\pm$ 0.015 -0.091 0.895 $\pm$ 0.011 -0.055 0.810 $\pm$ 0.013 -0.055
+w/o Transformer 0.885 $\pm$ 0.012 -0.075 0.620 $\pm$ 0.015 -0.101 0.870 $\pm$ 0.013 -0.080 0.765 $\pm$ 0.016 -0.100
+Temporal only 0.900 $\pm$ 0.011 -0.060 0.640 $\pm$ 0.013 -0.081 0.905 $\pm$ 0.010 -0.045 0.820 $\pm$ 0.012 -0.045
+Feature only 0.915 $\pm$ 0.009 -0.045 0.660 $\pm$ 0.014 -0.061 0.865 $\pm$ 0.012 -0.085 0.760 $\pm$ 0.015 -0.105
+w/ LayerNorm 0.940 $\pm$ 0.010 -0.020 0.695 $\pm$ 0.012 -0.026 0.925 $\pm$ 0.011 -0.025 0.835 $\pm$ 0.013 -0.030
+Transformer Only 0.875 $\pm$ 0.013 -0.085 0.605 $\pm$ 0.016 -0.116 0.845 $\pm$ 0.014 -0.105 0.730 $\pm$ 0.017 -0.135
+Table 23: Ablation study results on CKD and MIMIC-IV datasets. Values denote mean $\pm$ 95% confidence intervals over 5 runs. $\Delta$ shows absolute performance drop from the full CoI model.
+Appendix E Chain-of-Influence Visualizations
+Figure 5: Chain-of-Influence network visualization for CKD progression prediction.
+Nodes represent clinical features at specific time points (e.g., eGFR_t-5 indicates eGFR at 5 time steps before prediction).
+Directed edges capture learned temporal dependencies between features, with red edges indicating risk-enhancing influences and blue edges representing protective relationships.
+The network reveals complex cascading pathways from early cardiovascular events through kidney function decline to healthcare utilization escalation, demonstrating CoI’s ability to capture clinically meaningful temporal-feature interactions that drive disease progression.

### v3

@@
 1 introduction
-Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens NKF_CKD_2024; NCHS2019Mortality.
-The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging lin2013progression; li2024enhancing.
+Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens National Kidney Foundation (2024); Kochanek et al.
+(2020).
+The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging Lin et al.
+(2013); Li and Padman (2026).
 Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities.
-However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care caruana2015intelligible; rudin2019stop; tonekaboni2019clinicians.
-This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME ribeiro2016should and SHAP lundberg2017unified toward more sophisticated attention-based approaches that provide intrinsic interpretability choi2016retain; choi2017gram; bardhan2024icu.
+However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care Caruana et al.
+(2015); Rudin (2019); Tonekaboni et al.
+(2019).
+This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME Ribeiro et al.
+(2016) and SHAP Lundberg and Lee (2017) toward more sophisticated attention-based approaches that provide intrinsic interpretability Choi et al.
+(2016; 2017); Ma et al.
+(2017).
 Despite these advances, current approaches to clinical prediction face fundamental limitations.
 Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships.
-Even sophisticated attention-based models like RETAIN choi2016retain, while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
+Even sophisticated attention-based models like RETAIN Choi et al.
+(2016), while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
 These chains of influence—where early clinical indicators propagate through interconnected pathways to affect later outcomes—represent a core challenge for prediction in both chronic and acute care settings.
@@
 Attention mechanisms marked a turning point for interpretability in clinical deep learning.
-The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning choi2016retain.
+The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning Choi et al.
+(2016).
 By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions.
-Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies ma2017dipole.
+Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies Ma et al.
+(2017).
 While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction.
@@
 To address this, a second wave of research developed more sophisticated temporal attention models.
-For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time zhang2019attain.
+For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time Zhang (2019).
 Other models introduced domain-specific structural priors.
-StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time gao2020stagenet.
-Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component ma2020adacare.
+StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time Gao et al.
+(2020).
+Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component Ma et al.
+(2020).
 These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
 Self-Attention and Transformer-Based Approaches
-The Transformer and its self-attention mechanism provide a powerful way to model global, all-to-all interactions within a sequence vaswani2017attention.
-In the clinical domain, BEHRT treats a patient record as a sequence of concepts and learns contextualized representations li2020behrt; PAVE builds on this with multi-layer self-attention to identify important events and aggregate them into higher-level clinical patterns steinberg2021language; and XTSFormer introduces multi-scale temporal attention to separate short-term acute signals from long-term chronic trends zhang2024xtsformer.
+The Transformer and its self-attention mechanism provide a powerful way to model global, all-to-all interactions within a sequence Vaswani et al.
+(2017).
+In the clinical domain, BEHRT treats a patient record as a sequence of concepts and learns contextualized representations Li et al.
+(2020); PAVE builds on this with multi-layer self-attention to identify important events and aggregate them into higher-level clinical patterns Steinberg et al.
+(2021); and XTSFormer introduces multi-scale temporal attention to separate short-term acute signals from long-term chronic trends Xiao et al.
+(2025).
 While self-attention captures a dense web of pairwise dependencies, the resulting maps remain difficult to interpret clinically: they highlight correlations but do not clearly expose directed, sequential pathways of influence.
@@
 3.1.2 Temporal and Feature-Level Attention
-Inspired by the two-level attention framework in RETAIN choi2016retain, we employ two separate bidirectional LSTM modules that operate in parallel on $\mathbf{E}_{\text{t}}$ to extract complementary temporal and feature-level importance scores.
+Inspired by the two-level attention framework in RETAIN Choi et al.
+(2016), we employ two separate bidirectional LSTM modules that operate in parallel on $\mathbf{E}_{\text{t}}$ to extract complementary temporal and feature-level importance scores.
 Unlike RETAIN’s use of GRU units, our bidirectional LSTMs capture long-range dependencies from both past and future contexts, which is critical for detecting subtle temporal-feature interactions in clinical sequences.
@@
 The feature-weighted embeddings $\widetilde{\mathbf{E}}$ are passed through $L$ stacked Transformer layers to model temporal dependencies and inter-feature interactions.
-Each layer $l$ applies multi-head self-attention ($N_{h}$ heads) and a feed-forward network, with residual connections and Dynamic Tanh (DyT) normalization zhu2025transformers after each sub-layer.
+Each layer $l$ applies multi-head self-attention ($N_{h}$ heads) and a feed-forward network, with residual connections and Dynamic Tanh (DyT) normalization Zhu et al.
+(2025) after each sub-layer.
 Given input $\mathbf{Z}^{(l-1)}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}$ (with $\mathbf{Z}^{(0)}=\widetilde{\mathbf{E}}$), the attention output is $\mathbf{U}^{(l)}$, and attention weights from all layers are retained for interpretability (Section 3.2).
@@
 4.1.2 MIMIC-IV Dataset
-To assess generalizability in fast-changing conditions, we use MIMIC-IV v3.1 johnson2023mimic and construct a cohort from first ICU stays (65,366 patients out of 94,458 stays).
+To assess generalizability in fast-changing conditions, we use MIMIC-IV v3.1 Johnson et al.
+(2023) and construct a cohort from first ICU stays (65,366 patients out of 94,458 stays).
 Raw ICU measurements are recorded at irregular times; for modeling, we aggregate events into 1-hour windows over the first 48 hours and retain irregularity signals via masks and time-gap features (time since last observation) across 15 variables (vitals, labs, demographics).
@@
 We evaluate model performance using standard classification metrics.
-The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) hanley1982meaning; fawcett2006introduction, which measures the model’s discriminative ability across all classification thresholds.
-We also report the F1-Score van1979information, the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets.
-Additionally, we assess overall prediction accuracy sokolova2009systematic, which captures the proportion of correctly classified instances across both classes.
-Finally, we examine precision van1979information, which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions powers2011evaluation.
+The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) Hanley and McNeil (1982); Fawcett (2006), which measures the model’s discriminative ability across all classification thresholds.
+We also report the F1-Score Powers (2020), the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets.
+Additionally, we assess overall prediction accuracy Sokolova and Lapalme (2009), which captures the proportion of correctly classified instances across both classes.
+Finally, we examine precision Powers (2020), which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions Powers (2020).
 4.4 Training and Implementation Details
-Before training, we impute missing values using a MICE-based technique raghunathan2001multivariate; van2007multiple applied separately on each dataset.
+Before training, we impute missing values using a MICE-based technique Raghunathan et al.
+(2001); Van Buuren (2007) applied separately on each dataset.
 Given the substantial class imbalance in both cohorts, we evaluate several mitigation strategies—including SMOTE, ADASYN, class-weighted loss, and Temporal SMOTE (TSMOTE)—and find that TSMOTE offers the best F1–AUROC trade-off (Detailed result is provided in Appendix C.2), so we adopt it for all main experiments.
@@
 CoI thus goes beyond generic recency bias and concentrates on clinically meaningful temporal windows where deterioration becomes effectively irreversible.
-At the feature level (Figure 2(b)2(c)), both models consistently identify key predictors, with diabetes emerging as the dominant risk factor, in line with clinical evidence burckhardt2017multi; li2025towards.
+At the feature level (Figure 2(b)2(c)), both models consistently identify key predictors, with diabetes emerging as the dominant risk factor, in line with clinical evidence Burckhardt et al.
+(2017); Li et al.
+(2025).
 Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels also rank among the top features.
@@
 The network reveals complex cascading pathways from early cardiovascular events through kidney function decline to healthcare utilization escalation, demonstrating CoI’s ability to capture clinically meaningful temporal-feature interactions that drive disease progression.
+Figure 6: Top-55 patient-specific Chain-of-Influence graph for a single patient, showing the most influential feature–time pairs and their directed paths into the prediction node.
+Time Interval: $t_{0}\rightarrow prediction$.
+Figure 7: Path-centric view anchored at CKD Stage 4 at $t_{3}$, highlighting the high-influence chain through subsequent visits that drives the final ESRD risk.